### v1

# Congruences modulo powers of 3 for 6-colored generalized Frobenius partitions

Abstract.
In 1984, Andrews introduced the family of partition functions $c\phi_{k}(n)$, the number of generalized Frobenius partitions of $n$ with $k$ colors. In 2016, Gu, Wang, and Xia proved some congruences about $c\phi_{6}(n)$ and gave a conjecture on congruences modulo powers of 3 for $c\phi_{6}(n)$. We solve the revised conjecture proposed by Gu, Wang, and Xia using a method similar to that of Banerjee and Smoot.
Key words and phrases:
Generalized Frobenius partitions; Congruences; Modular forms; Partitions
2010 Mathematics Subject Classification:
11P83, 05A17
1. Introduction
A partition of a positive integer $n$ is defined as a sequence of positive integers in non-increasing order that sum to $n$. Denoted by $p(n)$, it represents the count of such partitions for $n$. The following distinguished congruences were discovered and proved by Ramanujan in [12]:
$p(5n+4)\equiv 0$ $\pmod{5},$
$p(7n+5)\equiv 0$ $\pmod{7},$
$p(11n+6)\equiv 0$ $\pmod{11},$
for every non-negative integer $n$. In fact, for $\alpha\geq 1$,
(1.1) $p(5^{\alpha}n+\delta_{5,\alpha})$ $\equiv 0\pmod{5^{\alpha}},$
(1.2) $p(7^{\alpha}n+\delta_{7,\alpha})$ $\equiv 0\pmod{7^{[\frac{\alpha+2}{2}]}},$
(1.3) $p(11^{\alpha}n+\delta_{11,\alpha})$ $\equiv 0\pmod{11^{\alpha}},$
where $\delta_{t,\alpha}$ is the reciprocal modulo $t^{\alpha}$ of 24. (1.1) and (1.2) were first proved by Watson [14] in 1938. For an elementary of (1.1) see Hirschhorn and Hunt [10] and for an elementary proof of (1.2) see Garvan [7]. (1.3) was proved by Atkin [3] in 1967.
In 1984 AMS Memoir [2], Andrews introduced the concept of generalized Frobenius partitions of $n$, which is a notation of the from
$\left(\begin{array}[]{cccc}a_{1}&a_{2}&\cdots&a_{r}\\ b_{1}&b_{2}&\cdots&b_{r}\end{array}\right)$
where non-negative integers $a_{i}$, $b_{i}$ with
$n=r+\sum_{i=1}^{r}a_{i}+\sum_{i=1}^{r}b_{i},$
and each row is of the same length and each is arranged in non-increasing order.
Andrews [2] also discussed a variant of generalized Frobenius partition which is called a $k$-colored generalized Frobenius partition. A $k$-colored generalized Frobenius partition is an array of the above from where the integer entries are taken from $k$ distinct copies of non-negative integers distinguished by color, the rows are ordered first according to size and then according to color with no two consecutive like entries in any row. For any positive integer $k$, let $c\phi_{k}(n)$ denotes the number of $k$-colored generalized Frobenius partitions of $n$. The generating function for $c\phi_{k}(n)$ is given by [2]
$\sum_{n=0}^{\infty}c\phi_{k}(n)q^{n}=\prod_{n=1}^{\infty}\frac{1}{(1-q^{n})^{k}}\sum_{m_{1},m_{2},\cdots,m_{k-1}=-\infty}^{\infty}q^{Q(m_{1},m_{2},\cdots,m_{k-1})},$
where
$Q(m_{1},m_{2},\cdots,m_{k-1})=\sum_{j=1}^{k-1}m_{j}^{2}+\sum_{1\leq i\textless j\leq k-1}m_{i}m_{j}.$
In this paper, we restrict our attention to generalized $6$-colored Frobenius partitions. Baruah and Sarmash [6, Theorem 2.1] represented the generating function of $c\phi_{6}(n)$ in terms of Ramanujan’s theta functions:
(1.4) $\sum_{n=0}^{\infty}c\phi_{6}(n)q^{n}=$ $\frac{1}{(q;q)_{\infty}^{6}}\{\varphi^{3}(q)\varphi(q^{2})\varphi(q^{6})+24\psi^{3}(q)\psi(2^{2})\psi(q^{3})$
$+4q^{2}\varphi^{3}(q)\psi(q^{4})\psi(q^{12})\},$
where
$\varphi(q)=(-q;q^{2})_{\infty}^{2}(q^{2};q^{2})_{\infty},$
$\psi(q)=(-q;q^{2})_{\infty}(q^{4};q^{4})_{\infty}.$
They [6] also conjectured that for $n\geq 0$,
(1.5) $c\phi_{6}(3n+2)\equiv 0\pmod{27}.$
In 2015, Xia [13] proved (1.5) by employing the generating function for $c\phi_{6}(3n+2)$ due to Baruah and Sarmah. Xia further conjectured that
$c\phi_{6}(9n+7)\equiv 0\pmod{27},$
which was proved by Hirschhorn in [11]. In 2016, Gu, Wang, and Xia [9] proved some congruences modulo powers of 3 for $c\phi_{6}(n)$ utilizing the generating function for $c\phi_{6}(3n+1)$ given by Hirschhorn [11]. For example, for $n\geq 0$,
$c\phi_{6}(19683n+11482)$ $\equiv 0\pmod{3^{7}},$
$c\phi_{6}(59049n+44287)$ $\equiv 0\pmod{3^{7}}.$
Besides, they gave the following conjecture [9, Conjecture 1.2]: for any integer $j\geq 8$, there exist positive integers $a$ and $k$ such that for all $n\geq 0$,
(1.6) $c\phi_{6}(3^{k}n+a)\equiv 0\pmod{3^{j}}.$
Considering about the conjecture above, we will prove the following theorem:
Theorem 1.1.
For integers $n\geq 0$, and $\alpha\geq 1$,
(1.7) $c\phi_{6}(3^{\alpha}n+\lambda_{\alpha})\equiv 0\pmod{3^{\lfloor\frac{\alpha}{2}\rfloor+2}},$
where
$\lambda_{2\alpha-1}=\frac{3^{2\alpha-1}+1}{4}~{}\text{and}~{}\lambda_{2\alpha}=\frac{3^{2\alpha+1}+1}{4}.$
In this paper, we prove this theorem by employing a similar method to that in [5], utilizing the recently developed localization method and internal algebraic structure. The paper is organized as follows. In Section 2 we introduce the necessary notations and definitions. In Section 3 we derive modular equations on $\Gamma_{0}(12)$. Section 4 is dedicated to the proof of Theorem 1.1.
2. Preliminaries
In this paper, we use the following conventions: =$\mathbb{N}=\{0,1,2,\cdots\}$ denotes the set of nonnegative integers. The complex upper half plane is denoted by :=$\mathbb{H}:=\{\tau\in\mathbb{C}:\Im(\tau)\textgreater 0\}$. We will also use the short hand notation for the Dedekind eta function:
$\eta_{n}(\tau):=\eta(n\tau),~{}~{}n\in\mathbb{Z},~{}\tau\in\mathbb{H}.$
Throughout this paper, for $x\in\mathbb{R}$ the symbol $\lfloor x\rfloor$ denotes the largest integer less than or equal to $x$, and $\lceil x\rceil$ denotes the smallest integer greater than or equal to $x$. For $n\in\mathbb{Z}$, $\pi(n)$ denotes the 3-adic order of n (i.e. the highest power of 3 that divides $n$).
Let $f=\sum_{n\in\mathbb{Z}}a_{n}q^{n}$, $f\neq 0$, be such that $a_{n}=0$ for almost all $n\textless 0$. We define the order of $f$ (with respect to $q$) as the smallest integer $N$ such that $a_{N}\neq 0$ and write $N=ord_{q}(f)$. For instance, we let $F=f\circ t=\sum_{n\in\mathbb{Z}}a_{n}t^{n}$ where $t=\sum_{n\geq 1}b_{n}q^{n}$, then the $t$-order of $F$ is defined to be the smallest integer $N$ such that $a_{N}\neq 0$ and write $N=ord_{t}(F)$.
Definition 2.1.
For $f:\mathbb{H}\rightarrow\mathbb{C}$ and $m\in\mathbb{N}^{\ast}$ we define $U_{m}(f):\mathbb{H}\rightarrow\mathbb{C}$ by
$U_{m}(f)(\tau):=\frac{1}{m}\sum_{\lambda=0}^{m-1}f(\frac{\tau+\lambda}{m}),~{}\tau\in\mathbb{H}.$
$U_{m}$ is linear (over $\mathbb{C}$); in addition, it is easy to verify that
$U_{mn}=U_{m}\circ U_{n}=U_{n}\circ U_{m},~{}m,n\in\mathbb{N}^{\ast}.$
The operators $U_{m}$, introduced by Atkin and Lehner [4], are closely related to Hecke operators. They typically arise in the context of partition congruences [1] mostly because of the property: if
$f(\tau)=\sum_{n=-\infty}^{\infty}f_{n}q^{n}~{}~{}(q=e^{2\pi i\tau}),$
then
$U_{m}(f)(\tau)=\sum_{n=-\infty}^{\infty}f_{mn}q^{n}.$
Definition 2.2.
For $f:\mathbb{H}\rightarrow\mathbb{C}$ we define $U_{A}(f),U_{B}(f):\mathbb{H}\rightarrow\mathbb{C}$ by $U_{A}(f):=U_{3}(Af)$ and $U_{B}(f):=U_{3}(Bf)$, where
$A:=\frac{\eta_{9}^{9}\eta_{4}^{2}\eta_{2}^{5}}{\eta_{36}^{2}\eta_{18}^{5}\eta_{1}^{9}}~{}~{}\text{and}~{}~{}B:=\frac{\eta_{9}\eta_{2}^{2}}{\eta_{18}^{2}\eta_{1}}.$
In the next proof, we use the following lemma to describe the powers of $y$ in $L_{\alpha}$.
Lemma 2.3.
For $f(1)=8$ and there exists a recurrence relation that for $s\geq 2$,
$f(s)=\left\{\begin{array}[]{ll}3f(s-1),{}&\text{if}~{}s~{}\text{is even,}\\ 3f(s-1)+8,{}&\text{if}~{}s~{}\text{is odd.}\end{array}\right.$
Then for $s\geq 1$, we have
$f(s)=\left\{\begin{array}[]{ll}3^{s+1}-3,{}&\text{if}~{}s~{}\text{is even,}\\ 3^{s+1}-1,{}&\text{if}~{}s~{}\text{is odd.}\end{array}\right.$
Definition 2.4.
We define the $U$-sequence $(L_{\alpha})_{\alpha\geq 0}$ by
$L_{0}:=\frac{\eta_{12}^{5}\eta_{3}\eta_{2}^{8}}{\eta_{24}^{2}\eta_{8}^{2}\eta_{6}^{4}\eta_{4}^{3}\eta_{1}^{3}}+24+4\frac{\eta_{24}^{2}\eta_{8}^{2}\eta_{3}\eta_{2}^{10}}{\eta_{12}\eta_{6}^{2}\eta_{4}^{9}\eta_{1}^{3}}$
and for $\alpha\geq 1$:
$L_{2\alpha-1}:=U_{A}(L_{2\alpha-2})~{}~{}~{}~{}~{}\text{and}~{}~{}~{}L_{2\alpha}:=U_{B}(L_{2\alpha-1}).$
The proof of the following lemma is completely analogous to [3, p. 23] and we omit it.
Lemma 2.5.
For $\alpha\in\mathbb{N}^{\ast}$, we have
$L_{2\alpha-1}=q^{-1}\prod_{n=1}^{\infty}\frac{(1-q^{n})(1-q^{3n})^{9}}{(1-q^{2n})^{2}(1-q^{6n})^{5}(1-q^{12n})^{2}}\sum_{n=0}^{\infty}c\phi_{6}(3^{2\alpha-1}n+\lambda_{2\alpha-1})q^{n},$
$L_{2\alpha}=\prod_{n=1}^{\infty}\frac{(1-q^{n})^{9}(1-q^{3n})}{(1-q^{2n})^{5}(1-q^{4n})^{2}(1-q^{6n})^{2}}\sum_{n=0}^{\infty}c\phi_{6}(2^{2\alpha}n+\lambda_{2\alpha})q^{n}.$
Definition 2.6.
Let $t,y,p_{0},p_{1}$ be functions defined on $\mathbb{H}$ as follows:
$t:=\frac{\eta_{12}^{4}\eta_{2}^{2}}{\eta_{6}^{2}\eta_{4}^{4}},~{}~{}~{}y:=\frac{\eta_{4}^{3}\eta_{3}}{\eta_{12}\eta_{1}^{3}},~{}~{}~{}p_{0}:=\frac{\eta_{12}^{4}\eta_{3}^{12}\eta_{2}^{8}}{\eta_{6}^{8}\eta_{4}^{12}\eta_{1}^{4}},~{}~{}~{}p_{1}:=\frac{\eta_{12}^{2}\eta_{3}^{6}\eta_{2}^{4}}{\eta_{6}^{4}\eta_{4}^{6}\eta_{1}^{2}},$
which have Laurent series expansions in powers of $q$ with coefficients in $\mathbb{Z}$.
3. Modular equation
Our proof of Gu, Wang, and Xia’s conjecture relies on the identities in the Appendix. All these identities can be proved using modular function theory. Regarding to the valence formula in the modular forms, Garvan has written a MAPLE package called ETA which can prove Group I, Group II, and Group III in the Appendix automatically. See
(3.1) $https://qseries.org/fgarvan/qmaple/ETA/$
The tutorial for this see [8]. As $L_{0}$ defined in Definition 2.4, $L_{0}=t^{-1}+27+3t+9t^{2}$ can be proved directly by the ETA MAPLE package. While, the identity of $L_{1}$ in the Appendix cannot be proved directly since $L_{0}$ is not an eta-quotient. However, we can use the ETA MAPLE package to prove that
(3.2) $U_{A}(t^{-1})=$ $y^{8}(3^{2}t^{-1}+217\cdot 3+839\cdot 3^{2}t+3140\cdot 3^{2}t^{2}+12070\cdot 3t^{3}-1118\cdot 3^{2}t^{4}$
$-2126\cdot 3^{3}t^{5}-380\cdot 3^{4}t^{6}+55\cdot 3^{5}t^{7}+19\cdot 3^{6}t^{8}+3^{7}t^{9});$
(3.3) $U_{A}(1)=$ $y^{8}(t^{-1}+62\cdot 3+100\cdot 3^{3}t+9712t^{2}+3442\cdot 3t^{3}-1588\cdot 3t^{4}-1588\cdot 3^{2}t^{5}$
$-64\cdot 3^{4}t^{6}+11\cdot 3^{5}t^{7}+2\cdot 3^{6}t^{8});$
(3.4) $U_{A}(t)=$ $y^{8}(47+328\cdot 3t+1100\cdot 3t^{2}+2792t^{3}-610\cdot 3t^{4}-344\cdot 3^{2}t^{5}-20\cdot 3^{3}t^{6}$
$+8\cdot 3^{3}t^{7}-3^{4}t^{8});$
(3.5) $U_{A}(t^{2})=$ $y^{8}(3^{2}+122\cdot 3t+121\cdot 3^{2}t^{2}+29\cdot 3^{3}t^{3}-274\cdot 3t^{4}-40\cdot 3^{2}t^{5}$
$+22\cdot 3^{3}t^{6}-14\cdot 3^{4}t^{7}-3^{5}t^{8}+2\cdot 3^{6}t^{9}-3^{5}t^{10}-3^{6}t^{11}).$
Since $L_{1}=U_{A}(L_{0})$ and $L_{0}=t^{-1}+27+3t+9t^{2}$, using (3.2)–(3.5) along with the identities $p_{0}=(1+t)^{4}$, we deduce that
(3.6) $L_{1}=$ $y^{8}p_{0}(4\cdot 3^{2}t^{-1}+71\cdot 3^{4}+2351\cdot 3^{3}t+89\cdot 3^{5}t^{2}-1975\cdot 3^{4}t^{3}+407\cdot 3^{5}t^{4}$
$-19\cdot 3^{7}t^{5}+11\cdot 3^{7}t^{6}-3^{8}t^{7}).$
Definition 3.1.
With $t=t(\tau)$ as in Definition 2.6 we define:
$a_{0}(t)=-t,~{}~{}~{}a_{1}(t)=3t^{2},~{}~{}~{}a_{2}(t)=-9t^{3}+6t.$
We define $s:\{0,1,2\}\times\{1,2,3\}\rightarrow\mathbb{Z}$ to be the unique function satisfying
(3.7) $a_{j}(t)=\sum_{l=1}^{3}s(j,l)3^{\lfloor\frac{2l+j}{3}\rfloor}t^{l}.$
Lemma 3.2.
For $\lambda\in\{0,1,2\}$, let
$t_{\lambda}(\tau):=t(\frac{\tau+\lambda}{3}),~{}~{}\tau\in\mathbb{H}.$
Then in the polynomial ring $\mathbb{C}(t)[X]$:
$X^{3}+a_{2}(t)X^{2}+a_{1}(t)X+a_{0}(t)=(X-t_{0})(X-t_{1})(X-t_{2}).$
Proof.
First we prove
$\prod_{\lambda=0}^{2}t_{\lambda}(\tau)=-a_{0}(t)=t.$
With $\omega:=e^{\frac{2\pi i}{3}}$, we have
$\prod_{\lambda=0}^{2}t_{\lambda}(\tau)$ $=\prod_{\lambda=0}^{2}q^{\frac{1}{3}}\omega^{\lambda}\prod_{n=1}^{\infty}\frac{(1-q^{4n})^{4}(1-\omega^{2\lambda n}q^{\frac{2n}{3}})^{2}}{(1-q^{2n})^{2}(1-\omega^{4\lambda n}q^{\frac{4n}{3}})^{4}}$
$=q\prod_{n=1}^{\infty}\prod_{\lambda=0}^{2}\frac{(1-q^{4n})^{4}(1-\omega^{2\lambda n}q^{\frac{2n}{3}})^{2}}{(1-q^{2n})^{2}(1-\omega^{4\lambda n}q^{\frac{4n}{3}})^{4}}$
$=q\prod_{n=1}^{\infty}\frac{(1-q^{4n})^{12}}{(1-q^{2n})^{6}}\prod_{n=1}^{\infty}\frac{(1-q^{2n})^{2}(1-q^{12n})^{4}}{(1-q^{6n})^{2}(1-q^{4n})^{4}}\prod_{n=1}^{\infty}\frac{(1-q^{2n})^{6}}{(1-q^{4n})^{12}}$
$=t,$
where using the fact that $\prod_{\lambda=0}^{2}(1-\omega^{\lambda n}z)$ equals $(1-z)^{3}$ if $3\mid n$, and $1-z^{3}$ otherwise. For $a_{1}(t)$,
$a_{1}(t)=$ $t_{0}(\tau)t_{1}(\tau)+t_{0}(\tau)t_{2}(\tau)+t_{1}(\tau)t_{2}(\tau)=3tU_{3}(t^{-1})=3t^{2},$
where the last equation was verified by the ETA MAPLE package. For $a_{2}(t)$,
$a_{2}(t)=$ $-t_{0}(\tau)-t_{1}(\tau)-t_{2}(\tau)=-3U_{3}(t)=-9t^{3}+6t,$
where the last equation was again verified by the ETA MAPLE package. ∎
Similarly, we can obtain the following definition and lemma without detailed proofs.
Definition 3.3.
With $y=y(\tau)$ as in Definition 2.6 we define:
$b_{0}(y)=-y^{3},~{}~{}~{}b_{1}(y)=8y^{3}-3y^{2}-3y+1,~{}~{}~{}b_{2}(y)=-16y^{3}+12y^{2}+3y-2.$
Lemma 3.4.
For $\lambda\in\{0,1,2\}$, let
$y_{\lambda}(\tau):=y(\frac{\tau+\lambda}{3}),~{}~{}\tau\in\mathbb{H}.$
Then in the polynomial ring $\mathbb{C}(y)[X]$:
$X^{3}+b_{2}(y)X^{2}+b_{1}(y)X+b_{0}(y)=(X-y_{0})(X-y_{1})(X-y_{2}).$
Lemma 3.5.
For $u:\mathbb{H}\rightarrow\mathbb{C}$, $j\in\mathbb{Z}$, and $l\in\mathbb{N}$ :
(3.8) $U_{3}(ut^{j+3})=-a_{0}(t)U_{3}(ut^{j})-a_{1}(t)U_{3}(ut^{j+1})-a_{2}(t)U_{3}(ut^{j+2});$
(3.9) $U_{3}(uy^{l+3})=-b_{0}(y)U_{3}(uy^{l})-b_{1}(y)U_{3}(uy^{l+1})-b_{2}(y)U_{3}(uy^{l+2}).$
Proof.
For $\lambda\in\{0,1,2\}$, Lemma 3.2 implies
$t_{\lambda}^{3}+a_{2}(t)t_{\lambda}^{2}+a_{1}(t)t_{\lambda}+a_{0}(t)=0.$
Multiplying both sides with $u_{\lambda}t_{\lambda}^{j}$ where $u_{\lambda}(\tau):=u((\tau+\lambda)/3)$ gives
$u_{\lambda}t_{\lambda}^{j+3}+a_{2}(t)u_{\lambda}t_{\lambda}^{j+2}+a_{1}(t)u_{\lambda}t_{\lambda}^{j+1}+a_{0}(t)u_{\lambda}t_{\lambda}^{j}=0.$
Summing both sides over all $\lambda$ from $\{0,1,2\}$ completes the proof of (3.8). We can prove (3.9) similarly. ∎
In the following proofs, we need to use both (3.8) and (3.9), while the coefficients of $y^{m}t^{k}$ are not easily described, where $m\in\mathbb{N}$ and $k\in\mathbb{Z}$. According to Group II in the Appendix, it is possible to assume that $U_{A}(p_{1}y^{m}t^{k})$ can be written in the form of $y^{3m+8}p_{0}\sum_{n}a_{n}t^{n}$, $a_{n}\in\mathbb{Z}$ by Lemma 4.3. Based on the identity $y=\frac{1}{1-3t}$, we can rewrite $b_{0}(y)$, $b_{1}(y)$, and $b_{2}(y)$ into
(3.10) $b_{0}(y)$ $=-y^{9}(1-2\cdot 3^{2}t+5\cdot 3^{3}t^{2}-20\cdot 3^{3}t^{3}+5\cdot 3^{5}t^{4}-2\cdot 3^{6}t^{5}+3^{6}t^{6});$
(3.11) $b_{1}(y)$ $=-y^{6}(-3+3^{2}t+3^{4}t^{2}-14\cdot 3^{3}t^{3}+3^{5}t^{4}+3^{6}t^{5}-3^{6}t^{6});$
(3.12) $b_{2}(y)$ $=-y^{3}(3+4\cdot 3^{2}t+3^{3}t^{2}-2\cdot 3^{3}t^{3}).$
Similarly, we have $U_{B}(p_{0}y^{m}t^{k})=y^{3m}p_{1}\sum_{n}a_{n}t^{n}$ in Lemma 4.3. According to Lemma 3.5, we can calculate $U_{A}(p_{1}y^{m}t^{k})$ and $U_{B}(p_{0}y^{m}t^{k})$ using the identities in Appendix for all $k\geq-1$ and $m\in\mathbb{N}$ recursively. For example,
$U_{A}(p_{1}t^{2})=$ $-\sum_{j=0}^{2}a_{j}(t)U_{A}(p_{1}t^{j-1})$
$=$ $y^{8}p_{0}(11+209\cdot 3t-22\cdot 3^{3}t^{2}-106\cdot 3^{3}t^{3}+259\cdot 3^{4}t^{4}+259\cdot 3^{4}t^{5}-11\cdot 3^{8}t^{6}$
$+68\cdot 3^{7}t^{7}-238\cdot 3^{6}t^{8}+7\cdot 3^{8}t^{9}+29\cdot 3^{8}t^{10}-16\cdot 3^{9}t^{11}+11\cdot 3^{9}t^{12}-3^{10}t^{13}),$
$U_{B}(p_{0}y^{3}t^{-1})=$ $-\sum_{l=0}^{2}b_{l}(y)U_{B}(p_{0}y^{l}t^{-1})$
$=$ $y^{9}p_{1}(37\cdot 3+2992\cdot 3t+13628\cdot 3^{2}t^{2}+3872\cdot 3^{4}t^{3}-4814\cdot 3^{4}t^{4}-7600\cdot 3^{4}t^{5}$
$+2564\cdot 3^{5}t^{6}-3^{10}t^{8}).$
4. Proof of Theorem 1.1
Lemma 4.1.
Let $v,u:\mathbb{H}\rightarrow\mathbb{C}$ and $l\in\mathbb{Z}$. Suppose for $l\leq k\leq l+2$, there exist polynomials $p_{k}(t)\in\mathbb{Z}[t,t^{-1}]$ such that
(4.1) $U_{3}(ut^{k})=vp_{k}(t)$
and
(4.2) $ord_{t}(p_{k}(t))\geq\lceil\frac{k+s}{3}\rceil,$
for a fixed integer $s$. Then there exist families of polynomials $p_{k}(t)\in\mathbb{Z}[t,t^{-1}]$, integer $k\in\mathbb{Z}$, such that (4.1) and (4.2) hold for all $k\in\mathbb{Z}$.
Proof.
Let $N\textgreater l+2$ be an integer and assume by induction that there are families of polynomials $p_{k}(t)$ such that (4.1) and (4.2) hold for $l\leq k\leq N-1$. Suppose
$p_{k}(t)=\sum_{n\geq\lceil\frac{k+s}{3}\rceil}c(k,n)t^{n},~{}~{}~{}l\leq k\leq N-1,$
with integers $c(k,n)$. Apply Lemma 3.5 we get:
$U_{3}(ut^{N})=$ $-\sum_{j=0}^{2}a_{j}(t)U_{3}(ut^{N+j-3})$
$=$ $-\sum_{j=0}^{2}a_{j}(t)v\sum_{n\geq\lceil\frac{N+j-3+s}{3}\rceil}c(N+j-3,n)t^{n}$
$=$ $-v\sum_{j=0}^{2}a_{j}(t)t^{-1}\sum_{n\geq\lceil\frac{N+j+s}{3}\rceil}c_{i}(N+j-3,n-1)t^{n}.$
Recalling the fact that $a_{j}(t)t^{-1}$ for $0\leq j\leq 2$ is a polynomial in $t$, this determines polynomials $p_{N}(t)$ with the desired properties. ∎
Lemma 4.2.
Let $v,u:\mathbb{H}\rightarrow\mathbb{C}$ and $l\in\mathbb{Z}$. Suppose for $l\leq k\leq l+2$ there exist polynomials $p_{k}(t)\in\mathbb{Z}[t,t^{-1}]$, such that
(4.3) $U_{3}(ut^{k})=vp_{k}(t)$
where
(4.4) $p_{k}(t)=\sum_{n}c(k,n)3^{\lfloor\frac{2n-k+\gamma}{3}\rfloor}t^{n},$
with integers $\gamma$ and $c(k,n)$. Then there exist families of polynomials $p_{k}(t)\in\mathbb{Z}[t,t^{-1}]$, $k\in\mathbb{Z}$, of the form (4.4) for which property (4.3) holds for all $k\in\mathbb{Z}$.
Proof.
Suppose for an integer $N\textgreater l+2$ there are families of polynomials $p_{k}(t)$, of the form (4.4) satisfying property (4.3) for $l\leq k\leq N-1$. We proceed by mathematical induction on $N$. Applying Lemma 3.5 and using the induction base (4.3) and (4.4) we attain:
$U_{3}(ut^{N})=-\sum_{j=0}^{2}a_{j}(t)v\sum_{n}c(N+j-3,n)3^{\lfloor{\frac{2n-(N+j-3)+\gamma}{3}}\rfloor}t^{n}.$
Utilizing (3.7) this rewrites into:
$U_{3}(ut^{N})$ $=-\sum_{j=0}^{2}\sum_{l=1}^{3}s(j,l)3^{\lfloor{\frac{2l+j}{3}}\rfloor}t^{l}v\sum_{n}c(N+j-3,n)3^{\lfloor{\frac{2n-(N+j-3)+\gamma}{3}}\rfloor}t^{n}$
$=-v\sum_{j=0}^{2}\sum_{l=1}^{3}\sum_{n}s(j,l)c(N+j-3,n-l)3^{\lfloor{\frac{2(n-l)-(N+j-3)+\gamma}{3}}\rfloor+\lfloor{\frac{2l+j}{3}}\rfloor}t^{n}.$
The induction step is completed by simplifying the exponent of $3$ as follows:
$\lfloor{\frac{2(n-l)-(N+j-3)+\gamma}{3}}\rfloor+\lfloor{\frac{2l+j}{3}}\rfloor$
$\geq\lfloor{\frac{2(n-l)-(N+j-3)+\gamma+2l+j-2}{3}}\rfloor$
$\geq\lfloor{\frac{2n-N+\gamma}{3}}\rfloor.$
The induction proof for $N\textless l$ works analogously. ∎
Lemma 4.3.
Let $v,u:\mathbb{H}\rightarrow\mathbb{C}$, $l\in\mathbb{Z}$, and $r\in\mathbb{N}$. Suppose for $l\leq k\leq l+2$ and $r\leq m\leq r+2$ there exist polynomials $p_{m,k}(t)\in\mathbb{Z}[t,t^{-1}]$ such that
(4.5) $U_{3}(uy^{m}t^{k})=vy^{3m+a}p_{m,k}(t)$
and
(4.6) $ord_{t}(p_{m,k}(t))\geq\lceil\frac{k+s}{3}\rceil,$
for some fixed integers $a$ and $s$. Then there exist families of polynomials $p_{m,k}(t)\in\mathbb{Z}[t,t^{-1}]$, integer $k\in\mathbb{Z}$, such that (4.5) and (4.6) hold for all $m\in\mathbb{Z^{\ast}}$ and $k\in\mathbb{Z}$.
Proof.
By Lemma 4.1, we can directly get that for $r\leq m\leq r+2$, (4.5) and (4.6) hold for all $k\in\mathbb{Z}$. For a fixed integer $k$, let $N\textgreater r+2$ be an integer and assume by induction that there are families of polynomials $p_{m,k}(t)$ such that (4.5) and (4.6) hold for $r\leq m\leq N-1$. Suppose
$p_{m,k}(t)=\sum_{n\geq\lceil\frac{k+s}{3}\rceil}c(k,m,n)t^{n},~{}~{}~{}r\leq m\leq N-1,$
with integers $c(k,m,n)$. Apply Lemma 3.5 we get:
$U_{3}(uy^{N}t^{k})=$ $-\sum_{j=0}^{2}b_{j}(y)U_{3}(uy^{N+j-3}t^{k})$
$=$ $-\sum_{j=0}^{2}b_{j}(y)vy^{3(N+j-3)+a}\sum_{n\geq\lceil\frac{k+s}{3}\rceil}c(k,m,n)t^{n}.$
Recalling the equations (3.10)–(3.12) this determines (4.5) and (4.6) hold for all $m\in\mathbb{N}$. Thus we complete the proof of this lemma. ∎
Lemma 4.4.
Let $v,u:\mathbb{H}\rightarrow\mathbb{C}$, $l\in\mathbb{Z}$, and $r\in\mathbb{N}$. Suppose for $l\leq k\leq l+2$ and $r\leq m\leq r+2$ there exist polynomials $p_{m,k}(t)\in\mathbb{Z}[t,t^{-1}]$ such that
(4.7) $U_{3}(uy^{m}t^{k})=vy^{3m+a}p_{m,k}(t)$
and
(4.8) $p_{m,k}(t)=\sum_{n}c(k,m,n)3^{\lfloor\frac{2n-k+\gamma}{3}\rfloor}t^{n},$
with integers $\gamma$ and $c(k,m,n)$. Then there exist families of polynomials $p_{m,k}(t)\in\mathbb{Z}[t,t^{-1}]$, $k\in\mathbb{Z}$, of the form (4.8) for which property (4.7) holds for all $m\in\mathbb{N}$ and $k\in\mathbb{Z}$.
Proof.
By Lemma 4.2, we can directly get that for $r\leq m\leq r+2$, (4.7) and (4.8) hold for all $k\in\mathbb{Z}$. For a fixed integer $k$, let $N\textgreater r+2$ be an integer and assume by induction that there are families of polynomials $p_{m,k}(t)$ such that (4.7) and (4.8) hold for $r\leq m\leq N-1$. We proceed by mathematical induction on N. Applying Lemma 3.5 and using the induction base (4.7) and (4.8) we attain:
$U_{3}(uy^{N}t^{k})=-\sum_{j=0}^{2}b_{j}(y)U_{3}(uy^{N+j-3}t^{k})=-v\sum_{j=0}^{2}b_{j}(y)y^{3(N+j-3)+a}\sum_{n}c(k,m,n)3^{\lfloor\frac{2n-k+\gamma}{3}\rfloor}t^{n}.$
Since $b_{j}(y)$ can be written into $y^{9-3j}f(t)$, where $j\in\{0,1,2\}$ and $ord_{t}(f(t))\geq 0$, then (4.7) and (4.8) hold for $m=N$. Thus we complete the proof of this lemma. ∎
Definition 4.5.
A map $a:\mathbb{Z}\rightarrow\mathbb{Z}$ is called a discrete function if it has finite support. A map $a:\mathbb{Z}\times\mathbb{N}\times\mathbb{Z}\rightarrow\mathbb{Z}$ is called discrete array if for each $i\in\mathbb{Z}$, $j\in\mathbb{N}$ the map $a(i,j,-):\mathbb{Z}\rightarrow\mathbb{Z},k\mapsto a(i,j,k)$, has finite support.
Lemma 4.6.
Given $U_{A}(f)$, $U_{B}(f)$ as in Definition 2.2, $p_{0},p_{1}$, and $y$ as in Definition 2.6. Then there exist discrete arrays $a$ and $b$, such that the following relations hold for all $m\in\mathbb{N}$, $k\in\mathbb{Z}$:
(4.9) $U_{A}(p_{1}y^{m}t^{k})=y^{3m+8}p_{0}\sum_{n\geq\lceil\frac{k-3}{3}\rceil}a(k,m,n)t^{n},\text{where}~{}\pi(a(k,m,n))\geq\small{\lfloor\frac{2n-k+3}{3}\rfloor};$
(4.10) $U_{B}(p_{0}y^{m}t^{k})=y^{3m}p_{1}\sum_{n\geq\lceil\frac{k-1}{3}\rceil}b(k,m,n)t^{n},\text{where}~{}\pi(b(k,m,n))\geq{\lfloor\frac{2n-k+2}{3}\rfloor}.$
Proof.
The Appendix lists eighteen fundamental relations. The nine fundamental relations of Group II fit the pattern of the relation (4.9) for three values of $m$ and three values of $k$. The same observations apply to the Groups III with regard to the relations (4.10). Applying Lemmas 4.14.4, we immediately prove the statement for all $m\in\mathbb{N}$ and $k\in\mathbb{Z}$. ∎
In the Appendix, we mention the identity that $y=\frac{1}{1-3t}$. In the following proof, we will express each $L_{\alpha}$ as a rational polynomial in which the denominator consists of powers of $1-3t$. With in this mind, we define the multiplicative closed set
$\mathcal{S}:=\{(1-3t)^{n}:n\in\mathbb{Z}_{n\geq 0}\}.$
We will be working with subspaces of the localized ring
$\mathbb{Z}[t,t^{-1}]_{\mathcal{S}}:=\mathcal{S}^{-1}\mathbb{Z}[t,t^{-1}].$
Given the form of $L_{1}$ in (3.6), we expect to represent each $L_{\alpha}$ in terms of $\frac{t^{n}}{(1-3t)^{m}}$ for $n\geq-1$ and $m\geq 0$. If we set
$L_{2\alpha-1}=p_{0}y^{3^{2\alpha}-1}\sum_{n\geq-1}d_{n}^{(2\alpha-1)}t^{n};$
$L_{2\alpha}=p_{1}y^{3^{2\alpha+1}-3}\sum_{n\geq 0}d_{n}^{(2\alpha)}t^{n},$
for $\alpha\geq 1$. Recalling that $\pi(n)$ denotes the 3-adic order of n (i.e. the highest power of 3 that divides $n$), then Theorem 1.1 follows from the next lemma.
Lemma 4.7.
For integer $\alpha\geq 1$, we have
(4.11) $\pi(d_{n}^{(2\alpha-1)})\geq\lfloor\frac{2n+5}{3}\rfloor+\alpha,~{}~{}\text{for $n\geq-1$},$
(4.12) $\pi(d_{n}^{(2\alpha)})\geq\lfloor\frac{2n+5}{3}\rfloor+\alpha,~{}~{}\text{for $n\geq 1$}.$
In particular,
(4.13) $\pi(d_{0}^{(2\alpha-1)})\geq\alpha+2,$
(4.14) $\pi(d_{0}^{(2\alpha)})\geq\alpha+2.$
Proof.
We proceed by induction on $\alpha$. From (3.6), we have
$L_{1}=$ $y^{8}p_{0}(4\cdot 3^{2}t^{-1}+71\cdot 3^{4}+2351\cdot 3^{3}t+89\cdot 3^{5}t^{2}-1975\cdot 3^{4}t^{3}+407\cdot 3^{5}t^{4}$
$-19\cdot 3^{7}t^{5}+11\cdot 3^{7}t^{6}-3^{8}t^{7}),$
which demonstrates that (4.11) and (4.13) hold for $\alpha=1$. Next, we suppose (4.11) and (4.13) hold for $\alpha\leq s$. Then by Definition 2.4 and Lemma 4.6, we have
$L_{2s}=$ $U_{B}(L_{2s-1})$
$=$ $\sum_{k\geq-1}d_{k}^{(2s-1)}U_{B}(p_{0}y^{3^{2s}-1}t^{k})$
$=$ $p_{1}y^{3^{2s+1}-3}\sum_{k\geq-1}d_{k}^{(2s-1)}\sum_{n\geq\lceil\frac{k-1}{3}\rceil}b(k,3^{2s}-1,n)t^{n}$
$=$ $p_{1}y^{3^{2s+1}-3}\sum_{n\geq 0}\sum_{-1\leq k\leq 3n+1}d_{k}^{(2s-1)}b(k,3^{2s}-1,n)t^{n}.$
By comparing the coefficients of $p_{1}y^{3^{2s+1}-3}t^{n}$ in the equation above, we have
(4.15) $d_{n}^{(2s)}=\sum_{-1\leq k\leq 3n+1}d_{k}^{(2s-1)}b(k,3^{2s}-1,n).$
Hence, for $n\geq 1$, we have
$\pi(d_{n}^{(2s)})$ $\geq\min_{k\geq-1}\{\lfloor\frac{2k+5}{3}\rfloor+s+\lfloor\frac{2n-k+2}{3}\rfloor\}$
$\geq\lfloor\frac{2n+5}{3}\rfloor+s,$
which indicates (4.12) holds for $\alpha=s$. For $d_{0}^{(2s)}$, from (4.15) we obtain that
$d_{0}^{(2s)}=$ $d_{-1}^{(2s-1)}b(-1,3^{2s}-1,0)+d_{0}^{(2s-1)}b(0,3^{2s}-1,0)+d_{1}^{(2s-1)}b(1,3^{2s}-1,0).$
By the assumption, we have $\pi(d_{-1}^{(2s-1)})\geq s+1$, $\pi(d_{0}^{(2s-1)})\geq s+2$, and $\pi(d_{1}^{(2s-1)})\geq s+2$. We denote $[t^{k}]f$ to the coefficients of $t^{k}$ in $f$. With (3.9), we have
(4.16) $b(-1,m,0)=$ $[p_{1}y^{3m}]U_{B}(p_{0}y^{m}t^{-1})$
$=$ $[p_{1}y^{3m}]\{y^{9}(1-2\cdot 3^{2}t+5\cdot 3^{3}t^{2}-20\cdot 3^{3}t^{3}+5\cdot 3^{5}t^{4}-2\cdot 3^{6}t^{5}+3^{6}t^{6})$
$\cdot U_{B}(p_{0}y^{m-3}t^{-1})+y^{6}(-3+3^{2}t+3^{4}t^{2}-14\cdot 3^{3}t^{4}+3^{6}t^{5}-3^{6}t^{6})$
$\cdot U_{B}(p_{0}y^{m-2}t^{-1})+y^{3}(3+4\cdot 3^{2}t+3^{3}t^{2}-2\cdot 3^{3}t^{3})U_{B}(p_{0}y^{m-1}t^{-1})\}$
$=$ $b(-1,m-3,0)-3b(-1,m-2,0)+3b(-1,m-1,0).$
In the Appendix, we get that $b(-1,0,0)=12$, $b(-1,1,0)=36$, and $b(-1,2,0)=69$ which all can be divided by $3$. By the recurrence relation above, we attain that $3\mid b(-1,m,0)$ for all $m\in\mathbb{N}$. Thus, we get that (4.14) holds for $\alpha=s$.
Finally, we suppose (4.12) and (4.14) hold for $\alpha\leq s$. Then by Definition 2.4 and Lemma 4.6, we have
$L_{2s+1}=$ $U_{A}(L_{2s})$
$=$ $\sum_{k\geq 0}d_{k}^{(2s)}U_{A}(p_{1}y^{3^{2s+1}-3}t^{k})$
$=$ $p_{0}y^{3^{2s+2}-1}\sum_{k\geq 0}d_{k}^{(2s)}\sum_{n\geq\lceil\frac{k-3}{3}\rceil}a(k,3^{2s+1}-3,n)t^{n}$
$=$ $p_{0}y^{3^{2s+2}-1}\sum_{n\geq-1}\sum_{0\leq k\leq 3n+3}d_{k}^{(2s)}a(k,3^{2s+1}-3,n)t^{n}.$
Equating the coefficients of $p_{0}y^{3^{2s+2}-1}t^{n}$ in the equation above, we have
(4.17) $d_{n}^{(2s+1)}=\sum_{0\leq k\leq 3n+3}d_{k}^{(2s)}a(k,3^{2s+1}-3,n).$
Hence, for $n\geq-1$, we have
$\pi(d_{n}^{(2s+1)})$ $\geq\min\{\min_{k\geq 1}\{\lfloor\frac{2k+5}{3}\rfloor+s+\lfloor\frac{2n-k+3}{3}\rfloor\},s+2+\lfloor\frac{2n+3}{3}\rfloor\}$
$\geq\lfloor\frac{2n+5}{3}\rfloor+s+1,$
which indicates (4.11) holds for $\alpha=s+1$. From (4.17), we have
$d_{0}^{(2s+1)}=$ $d_{0}^{(2s)}a(0,3^{2s+1}-3,0)+d_{1}^{(2s)}a(1,3^{2s+1}-3,0)+d_{2}^{(2s)}a(2,3^{2s+1}-3,0)$
$+d_{3}^{(2s)}a(3,3^{2s+1}-3,0).$
By the assumption, we have $\pi(d_{0}^{(2s)})\geq s+2$, $\pi(d_{1}^{(2s)})\geq s+2$, $\pi(d_{2}^{(2s)})\geq s+3$, and $\pi(d_{3}^{(2s)})\geq s+3$. Similar to (4.16), we obtain that
$a(0,m,0)=$ $a(0,m-3,0)-18a(0,m-3,-1)-3a(0,m-2,0)+9a(0,m-2,-1)$
$+3a(0,m-1,0)+36a(0,m-1,-1),$
$a(1,m,0)=$ $a(1,m-3,0)-3a(1,m-2,0)+3a(1,m-1,0).$
From the Appendix, we get that
$a(0,0,0)=285,\hskip 10.0pta(0,1,0)=600,\hskip 10.0pta(0,2,0)=1068;$
$a(1,0,0)=66,\hskip 16.0pta(1,1,0)=108,\hskip 10.0pta(1,2,0)=159.$
Hence, we attain that $3\mid a(0,m,0)$ and $3\mid a(1,m,0)$ for all $m\in\mathbb{N}$ so that (4.13) holds for $\alpha=s+1$, which completes the proof of Lemma 4.7. ∎
Appendix. The fundamental relations
Group I:
$y=\frac{1}{1-3t},\hskip 10.0ptp_{0}=(1+t)^{4},\hskip 10.0ptp_{1}=(1+t)^{2},\hskip 10.0ptL_{0}=t^{-1}+3^{3}+3t+3^{2}t^{2}.$
Group II:
$U_{A}(p_{1}t^{-1})=$ $y^{8}p_{0}(11t^{-1}+38\cdot 3^{3}+1085\cdot 3^{2}t+212\cdot 3^{3}t^{2}-961\cdot 3^{3}t^{3}+98\cdot 3^{4}t^{4}+3^{7}t^{5});$
$U_{A}(p_{1})=$ $y^{8}p_{0}(t^{-1}+95\cdot 3+16\cdot 3^{5}t+5\cdot 3^{3}t^{2}-308\cdot 3^{3}t^{3}+91\cdot 3^{4}t^{4}-20\cdot 3^{5}t^{5}+11\cdot 3^{5}t^{6}-3^{6}t^{7});$
$U_{A}(p_{1}t)=$ $y^{8}p_{0}(22\cdot 3+176\cdot 3^{2}t-11\cdot 3^{4}t^{2}-19\cdot 3^{4}t^{3}+8\cdot 3^{4}t^{4}+79\cdot 3^{4}t^{5}-88\cdot 3^{5}t^{6}+17\cdot 3^{7}t^{7}$
$-2\cdot 3^{9}t^{8}+11\cdot 3^{7}t^{9}-3^{8}t^{10});$
$U_{A}(p_{1}yt^{-1})=$ $y^{11}p_{0}(14t^{-1}+101\cdot 3^{3}+7760\cdot 3^{2}t+188\cdot 3^{7}t^{2}+9572\cdot 3^{3}t^{3}-16102\cdot 3^{4}t^{4}-128\cdot 3^{7}t^{5}$
$+20\cdot 3^{10}t^{6}-14\cdot 3^{9}t^{7}-3^{10}t^{8});$
$U_{A}(p_{1}y)=$ $y^{11}p_{0}(t^{-1}+200\cdot 3+284\cdot 3^{4}t+5752\cdot 3^{3}t^{2}+1234\cdot 3^{3}t^{3}-5912\cdot 3^{4}t^{4}+884\cdot 3^{5}t^{5}$
$+392\cdot 3^{5}t^{6}-55\cdot 3^{6}t^{7});$
$U_{A}(p_{1}yt)=$ $y^{11}p_{0}(4\cdot 3^{3}+802\cdot 3^{2}t+751\cdot 3^{4}t^{2}-224\cdot 3^{4}t^{3}-1892\cdot 3^{4}t^{4}+2116\cdot 3^{4}t^{5}$
$-514\cdot 3^{5}t^{6}+16\cdot 3^{8}t^{7}-32\cdot 3^{7}t^{8}+14\cdot 3^{7}t^{9}-3^{8}t^{10});$
$U_{A}(p_{1}y^{2}t^{-1})=$ $y^{14}p_{0}(17t^{-1}+215\cdot 3^{3}+31742\cdot 3^{2}t+148039\cdot 3^{3}t^{2}+657251\cdot 3^{3}t^{3}+119546\cdot 3^{4}t^{4}$
$-29324\cdot 3^{7}t^{5}-1930\cdot 3^{9}t^{6}+4745\cdot 3^{9}t^{7}+241\cdot 3^{10}t^{8}-10\cdot 3^{14}t^{9}+17\cdot 3^{12}t^{10}+3^{13}t^{11});$
$U_{A}(p_{1}y^{2})=$ $y^{14}p_{0}(t^{-1}+356\cdot 3+2963\cdot 3^{3}t+17042\cdot 3^{4}t^{2}+239206\cdot 3^{3}t^{3}+16864\cdot 3^{4}t^{4}$
$-32678\cdot 3^{6}t^{5}+692\cdot 3^{5}t^{6}+35045\cdot 3^{6}t^{7}-52\cdot 3^{11}t^{8}-17\cdot 3^{11}t^{9}+2\cdot 3^{12}t^{10});$
$U_{A}(p_{1}y^{2}t)=$ $y^{14}p_{0}(53\cdot 3+770\cdot 3^{3}t+5827\cdot 3^{4}t^{2}+29704\cdot 3^{4}t^{3}-1958\cdot 3^{5}t^{4}$
$-103100\cdot 3^{4}t^{5}+21722\cdot 3^{5}t^{6}+392\cdot 3^{8}t^{7}-967\cdot 3^{7}t^{8}+98\cdot 3^{7}t^{9}-3^{8}t^{10}).$
Group III:
$U_{B}(p_{0}t^{-1})=$ $y^{0}p_{1}(4\cdot 3-14\cdot 3t+11\cdot 3^{2}t^{2}-3^{4}t^{3}-2\cdot 3^{4}t^{4}+5\cdot 3^{4}t^{5}-3^{5}t^{6});$
$U_{B}(p_{0})=$ $y^{0}p_{1}(5-11\cdot 3t+2\cdot 3^{4}t^{2}-14\cdot 3^{3}t^{3}-3^{3}t^{4}+8\cdot 3^{5}t^{5}-13\cdot 3^{5}t^{6}+5\cdot 3^{6}t^{8}-3^{7}t^{9});$
$U_{B}(p_{0}t)=$ $y^{0}p_{1}(1-3^{2}t+14\cdot 3^{2}t^{2}-28\cdot 3^{3}t^{3}+61\cdot 3^{3}t^{4}+23\cdot 3^{4}t^{5}-20\cdot 3^{6}t^{6}+8\cdot 3^{7}t^{7}$
$+25\cdot 3^{6}t^{8}-23\cdot 3^{7}t^{9}+2\cdot 3^{8}t^{10}+5\cdot 3^{8}t^{11}-3^{9}t^{12});$
$U_{B}(p_{0}yt^{-1})=$ $y^{3}p_{1}(4\cdot 3^{2}+88\cdot 3t-55\cdot 3^{2}t^{2}+8\cdot 3^{4}t^{3}-10\cdot 3^{4}t^{4}+8\cdot 3^{4}t^{5}-3^{5}t^{6});$
$U_{B}(p_{0}y)=$ $y^{3}p_{1}(8+46\cdot 3t-16\cdot 3^{3}t^{2}+38\cdot 3^{3}t^{3}-64\cdot 3^{3}t^{4}+5\cdot 3^{5}t^{5}+8\cdot 3^{5}t^{6}-8\cdot 3^{6}t^{7}$
$+8\cdot 3^{6}t^{8}-3^{7}t^{9});$
$U_{B}(p_{0}yt)=$ $y^{3}p_{1}(1+8\cdot 3^{2}t-38\cdot 3^{2}t^{2}+16\cdot 3^{4}t^{3}-128\cdot 3^{3}t^{4}+56\cdot 3^{4}t^{5}+2\cdot 3^{7}t^{6}-40\cdot 3^{6}t^{7}$
$+64\cdot 3^{6}t^{8}-8\cdot 3^{7}t^{9}-2\cdot 3^{9}t^{10}+8\cdot 3^{8}t^{11}-3^{9}t^{12});$
$U_{B}(p_{0}y^{2}t^{-1})=$ $y^{6}p_{1}(23\cdot 3+802\cdot 3t+923\cdot 3^{2}t^{2}-140\cdot 3^{4}t^{3}-7\cdot 3^{5}t^{4}+38\cdot 3^{4}t^{5}-3^{5}t^{6});$
$U_{B}(p_{0}y^{2})=$ $y^{6}p_{1}(11+274\cdot 3t+14\cdot 3^{5}t^{2}-31\cdot 3^{5}t^{3}+314\cdot 3^{3}t^{4}-49\cdot 3^{5}t^{5}+62\cdot 3^{5}t^{6}-19\cdot 3^{6}t^{7}$
$+11\cdot 3^{6}t^{8}-3^{7}t^{9});$
$U_{B}(p_{0}y^{2}t)=$ $y^{6}p_{1}(1+28\cdot 3^{2}t+20\cdot 3^{4}t^{2}-197\cdot 3^{3}t^{3}+430\cdot 3^{3}t^{4}-280\cdot 3^{4}t^{5}+5\cdot 3^{8}t^{6}-32\cdot 3^{6}t^{7}$
$-32\cdot 3^{6}t^{8}+40\cdot 3^{7}t^{9}-17\cdot 3^{8}t^{10}+11\cdot 3^{8}t^{11}-3^{9}t^{12}).$
Acknowledgements
The first author was supported by the National Key R&D Program of China (Grant No. 2024YFA1014500). The first author was also supported in part by the National Natural Science Foundation of China (Grant No. 12201387).

### v2

@@
 Abstract.
-In 1984, Andrews introduced the family of partition functions $c\phi_{k}(n)$, the number of generalized Frobenius partitions of $n$ with $k$ colors.
-In 2016, Gu, Wang, and Xia proved some congruences about $c\phi_{6}(n)$ and gave a conjecture on congruences modulo powers of 3 for $c\phi_{6}(n)$.
-We solve the revised conjecture proposed by Gu, Wang, and Xia using a method similar to that of Banerjee and Smoot.
+In $1984$, Andrews introduced the family of partition functions $c\phi_{k}(n)$, which enumerate generalized Frobenius partitions of $n$ with $k$ colors.
+In $2016$, Gu, Wang, and Xia established several congruences for $c\phi_{6}(n)$ and proposed a conjecture concerning congruences modulo powers of $3$ for this function.
+In this paper, we resolve a revised version of their conjecture by employing an approach analogous to that developed by Banerjee and Smoot.
 Key words and phrases:
@@
 1. Introduction
-A partition of a positive integer $n$ is defined as a sequence of positive integers in non-increasing order that sum to $n$.
-Denoted by $p(n)$, it represents the count of such partitions for $n$.
-The following distinguished congruences were discovered and proved by Ramanujan in [12]:
+A partition of a positive integer $n$ is a non-increasing sequence of positive integers that sum to $n$.
+The number of such partitions is denoted by $p(n)$.
+Among the most celebrated results in partition theory are the following congruences, discovered and proved by Ramanujan [12]:
 $p(5n+4)\equiv 0$ $\pmod{5}$,
@@
 $p(11n+6)\equiv 0$ $\pmod{11}$,
-for every non-negative integer $n$. In fact, for $\alpha\geq 1$,
+for all non-negative integers $n$. In fact, for any integer $\alpha\geq 1$,
 (1.1) $p(5^{\alpha}n+\delta_{5,\alpha})$ $\equiv 0\pmod{5^{\alpha}}$,
@@
 (1.3) $p(11^{\alpha}n+\delta_{11,\alpha})$ $\equiv 0\pmod{11^{\alpha}}$,
-where $\delta_{t,\alpha}$ is the reciprocal modulo $t^{\alpha}$ of 24.
+where $\delta_{t,\alpha}$ is defined by the congruence $24\delta_{t,\alpha}\equiv 1\pmod{t^{\alpha}}$.
 (1.1) and (1.2) were first proved by Watson [15] in 1938.
-For an elementary of (1.1) see Hirschhorn and Hunt [10] and for an elementary proof of (1.2) see Garvan [7].
-(1.3) was proved by Atkin [3] in 1967.
-In 1984 AMS Memoir [2], Andrews introduced the concept of generalized Frobenius partitions of $n$, which is a notation of the from
+Later, elementary proofs of (1.1) and (1.2) were given by Hirschhorn and Hunt [10] and Garvan [7], respectively.
+The case for $11$, (1.3) was established by Atkin [3] in 1967.
+In his 1984 AMS Memoir [2], Andrews introduced the concept of generalized Frobenius partitions of $n$. It is defined as a two-rowed array of the form
 $\left(\begin{array}[]{cccc}a_{1}&a_{2}&\cdots&a_{r}\\ b_{1}&b_{2}&\cdots&b_{r}\end{array}\right)$
-where non-negative integers $a_{i}$, $b_{i}$ with
-$n=r+\sum_{i=1}^{r}a_{i}+\sum_{i=1}^{r}b_{i},$
-and each row is of the same length and each is arranged in non-increasing order.
-Andrews [2] also discussed a variant of generalized Frobenius partition which is called a $k$-colored generalized Frobenius partition.
-A $k$-colored generalized Frobenius partition is an array of the above from where the integer entries are taken from $k$ distinct copies of non-negative integers distinguished by color, the rows are ordered first according to size and then according to color with no two consecutive like entries in any row.
-For any positive integer $k$, let $c\phi_{k}(n)$ denotes the number of $k$-colored generalized Frobenius partitions of $n$.
-The generating function for $c\phi_{k}(n)$ is given by [2]
+where the entries $a_{i}$, $b_{i}$ are non-negative integers satisfying the condition
+$n=r+\sum_{i=1}^{r}a_{i}+\sum_{i=1}^{r}b_{i}.$
+Furthermore, both sequences $(a_{1},a_{2},\cdots,a_{r})$ and $(b_{1},b_{2},\cdots,b_{r})$ must be non-increasing.
+Andrews [2] also introduced a variant known as the $k$-colored generalized Frobenius partition.
+It is an array of the form described above, but whose entries are drawn from $k$ distinct copies of the non-negative integers, each copy distinguished by a color.
+The entries in each row are ordered first by their numerical value and subsequently by their color, with the additional constraint that no two consecutive entries in any row are identical in both value and color.
+For any positive integer $k$, let $c\phi_{k}(n)$ denote the number of such $k$-colored generalized Frobenius partitions of $n$. Andrews [2] established the following generating function:
 $\sum_{n=0}^{\infty}c\phi_{k}(n)q^{n}=\prod_{n=1}^{\infty}\frac{1}{(1-q^{n})^{k}}\sum_{m_{1},m_{2},\cdots,m_{k-1}=-\infty}^{\infty}q^{Q(m_{1},m_{2},\cdots,m_{k-1})},$
@@
 $Q(m_{1},m_{2},\cdots,m_{k-1})=\sum_{j=1}^{k-1}m_{j}^{2}+\sum_{1\leq i\textless j\leq k-1}m_{i}m_{j}.$
-In this paper, we restrict our attention to generalized $6$-colored Frobenius partitions.
-Baruah and Sarmash [6, Theorem 2.1] represented the generating function of $c\phi_{6}(n)$ in terms of Ramanujan’s theta functions:
-(1.4) $\sum_{n=0}^{\infty}c\phi_{6}(n)q^{n}=$ $\frac{1}{(q;q)_{\infty}^{6}}\{\varphi^{3}(q)\varphi(q^{2})\varphi(q^{6})+24\psi^{3}(q)\psi(2^{2})\psi(q^{3})$
+This paper focuses on generalized $6$-colored Frobenius partitions.
+Baruah and Sarmash [6, Theorem 2.1] expressed the generating function for $c\phi_{6}(n)$ in terms of Ramanujan’s theta functions as follows:
+(1.4) $\sum_{n=0}^{\infty}c\phi_{6}(n)q^{n}=$ $\frac{1}{(q;q)_{\infty}^{6}}\{\varphi^{3}(q)\varphi(q^{2})\varphi(q^{6})+24q\psi^{3}(q)\psi(q^{2})\psi(q^{3})$
 $+4q^{2}\varphi^{3}(q)\psi(q^{4})\psi(q^{12})\},$
@@
 $\psi(q)=(-q;q^{2})_{\infty}(q^{4};q^{4})_{\infty}.$
-They [6] also conjectured that for $n\geq 0$,
+They [6] also conjectured that for all $n\geq 0$,
 (1.5) $c\phi_{6}(3n+2)\equiv 0\pmod{27}$.
-In 2015, Xia [13] proved (1.5) by employing the generating function for $c\phi_{6}(3n+2)$ due to Baruah and Sarmah. Xia further conjectured that
+This conjecture was proved by Xia [14] using the generating function for $c\phi_{6}(3n+2)$ given by Baruah and Sarmah. Xia further conjectured that
 $c\phi_{6}(9n+7)\equiv 0\pmod{27}$,
-which was proved by Hirschhorn in [11].
-In 2016, Gu, Wang, and Xia [9] proved some congruences modulo powers of 3 for $c\phi_{6}(n)$ utilizing the generating function for $c\phi_{6}(3n+1)$ given by Hirschhorn [11].
-For example, for $n\geq 0$,
+which was subsequently established by Hirschhorn [11].
+In 2016, Gu, Wang, and Xia [9] proved several congruences modulo powers of 3 for $c\phi_{6}(n)$ by employing the generating function for $c\phi_{6}(3n+1)$ provided by Hirschhorn [11].
+For example, they showed that for all $n\geq 0$,
 $c\phi_{6}(19683n+11482)$ $\equiv 0\pmod{3^{7}}$,
 $c\phi_{6}(59049n+44287)$ $\equiv 0\pmod{3^{7}}$.
-Besides, they gave the following conjecture [9, Conjecture 1.2]: for any integer $j\geq 8$, there exist positive integers $a$ and $k$ such that for all $n\geq 0$,
+Additonally, they proposed the following conjecture [9, Conjecture 1.2]: for any integer $j\geq 8$, there exist positive integers $a$ and $k$ such that for all $n\geq 0$,
 (1.6) $c\phi_{6}(3^{k}n+a)\equiv 0\pmod{3^{j}}$.
-Considering about the conjecture above, we will prove the following theorem:
+Tang [13, Conjecture 1.4] confirmed the conjecture in the following explicit form, though without providing a proof:
 Theorem 1.1.
@@
 $\lambda_{2\alpha-1}=\frac{3^{2\alpha-1}+1}{4}~\text{and}~\lambda_{2\alpha}=\frac{3^{2\alpha+1}+1}{4}.$
-In this paper, we prove this theorem by employing a similar method to that in [5], utilizing the recently developed localization method and internal algebraic structure.
+Our proof follows an approach similar to Banerjee and Smoot [5], utilizing the recently developed localization method and internal algebraic structure.
 The paper is organized as follows.
@@
 In Section 3 we derive modular equations on $\Gamma_{0}(12)$.
-Section 4 is dedicated to the proof of Theorem 1.1.
+Section 4 presents the proof of Theorem 1.1.
 2. Preliminaries
-In this paper, we use the following conventions: =$\mathbb{N}=\{0,1,2,\cdots\}$ denotes the set of nonnegative integers.
-The complex upper half plane is denoted by :=$\mathbb{H}:=\{\tau\in\mathbb{C}:\Im(\tau)\textgreater 0\}$.
-We will also use the short hand notation for the Dedekind eta function:
-$\eta_{n}(\tau):=\eta(n\tau),~{}~{}n\in\mathbb{Z},~{}\tau\in\mathbb{H}.$
-Throughout this paper, for $x\in\mathbb{R}$ the symbol $\lfloor x\rfloor$ denotes the largest integer less than or equal to $x$, and $\lceil x\rceil$ denotes the smallest integer greater than or equal to $x$.
+In this paper, we adhere to the following conventions.
+The sets of integers, nonegative integers, and positive integers are denoted by $\mathbb{Z}$, $\mathbb{N}=\{0,1,2,\cdots\}$, and $\mathbb{N^{\ast}}=\{1,2,\cdots\}$, respectively.
+Let $\mathbb{H}:=\{\tau\in\mathbb{C}:\Im(\tau)\textgreater 0\}$ represent the complex upper half plane.
+We use the notation $\eta_{n}(\tau):=\eta(n\tau),~~n\in\mathbb{Z},~\tau\in\mathbb{H}$.
+For $\in\mathbb{R}, x\lfloor x\rfloor$ denotes the largest integer less than or equal to $x$, and $\lceil x\rceil$ denotes the smallest integer greater than or equal to $x$.
 For $n\in\mathbb{Z}$, $\pi(n)$ denotes the 3-adic order of n (i.e.
 the highest power of 3 that divides $n$).
-Let $f=\sum_{n\in\mathbb{Z}}a_{n}q^{n}$, $f\neq 0$, be such that $a_{n}=0$ for almost all $n\textless 0$.
-We define the order of $f$ (with respect to $q$) as the smallest integer $N$ such that $a_{N}\neq 0$ and write $N=ord_{q}(f)$.
-For instance, we let $F=f\circ t=\sum_{n\in\mathbb{Z}}a_{n}t^{n}$ where $t=\sum_{n\geq 1}b_{n}q^{n}$, then the $t$-order of $F$ is defined to be the smallest integer $N$ such that $a_{N}\neq 0$ and write $N=ord_{t}(F)$.
+Let $f=\sum_{n\in\mathbb{Z}}a_{n}q^{n}$ (with $f\neq 0$) be a series such that $a_{n}=0$ for all but finitely many $n\textless 0$.
+We define the order of $f$ (with respect to $q$) as the smallest integer $N$ for which $a_{N}\neq 0$, and write $ord_{q}(f)=N$.
+Furthermore, if $F=f\circ t=\sum_{n\in\mathbb{Z}}a_{n}t^{n}$, where $t=\sum_{n\geq 1}b_{n}q^{n}$, then the $t$-order of $F$ is defined as the smallest integer $N$ such that $a_{N}\neq 0$, denoted $N=ord_{t}(F)$.
 Definition 2.1.
@@
 $A:=\frac{\eta_{9}^{9}\eta_{4}^{2}\eta_{2}^{5}}{\eta_{36}^{2}\eta_{18}^{5}\eta_{1}^{9}}~~\text{and}~~B:=\frac{\eta_{9}\eta_{2}^{2}}{\eta_{18}^{2}\eta_{1}}.$
-In the next proof, we use the following lemma to describe the powers of $y$ in $L_{\alpha}$.
-Lemma 2.3.
-For $f(1)=8$ and there exists a recurrence relation that for $s\geq 2$,
-$f(s)=\left\{\begin{array}[]{ll}3f(s-1),{}&\text{if}~{}s~{}\text{is even,}\\ 3f(s-1)+8,{}&\text{if}~{}s~{}\text{is odd.}\end{array}\right.$
-Then for $s\geq 1$, we have
-$f(s)=\left\{\begin{array}[]{ll}3^{s+1}-3,{}&\text{if}~{}s~{}\text{is even,}\\ 3^{s+1}-1,{}&\text{if}~{}s~{}\text{is odd.}\end{array}\right.$
 Definition 2.3.
@@
 $L_{2\alpha-1}=q^{-1}\prod_{n=1}^{\infty}\frac{(1-q^{n})(1-q^{3n})^{9}}{(1-q^{2n})^{2}(1-q^{6n})^{5}(1-q^{12n})^{2}}\sum_{n=0}^{\infty}c\phi_{6}(3^{2\alpha-1}n+\lambda_{2\alpha-1})q^{n},$
-$L_{2\alpha}=\prod_{n=1}^{\infty}\frac{(1-q^{n})^{9}(1-q^{3n})}{(1-q^{2n})^{5}(1-q^{4n})^{2}(1-q^{6n})^{2}}\sum_{n=0}^{\infty}c\phi_{6}(2^{2\alpha}n+\lambda_{2\alpha})q^{n}.$
+$L_{2\alpha}=\prod_{n=1}^{\infty}\frac{(1-q^{n})^{9}(1-q^{3n})}{(1-q^{2n})^{5}(1-q^{4n})^{2}(1-q^{6n})^{2}}\sum_{n=0}^{\infty}c\phi_{6}(3^{2\alpha}n+\lambda_{2\alpha})q^{n}.$
 Definition 2.5.
@@
 3. Modular equation
-Our proof of Gu, Wang, and Xia’s conjecture relies on the identities in the Appendix.
-All these identities can be proved using modular function theory.
-Regarding to the valence formula in the modular forms, Garvan has written a MAPLE package called ETA which can prove Group I, Group II, and Group III in the Appendix automatically.
-See
+Our proof of the Gu-Wang-Xia conjecture relies on the modular identities in the Appendix.
+Many of these, specifically those in Groups II-IIIIII, can be automatically verified using Garvan’s MAPLE package ETA (see (3.1) and [8])
 (3.1) $https://qseries.org/fgarvan/qmaple/ETA/$
-The tutorial for this see [8].
-As $L_{0}$ defined in Definition 2.4, $L_{0}=t^{-1}+27+3t+9t^{2}$ can be proved directly by the ETA MAPLE package.
-While, the identity of $L_{1}$ in the Appendix cannot be proved directly since $L_{0}$ is not an eta-quotient.
-However, we can use the ETA MAPLE package to prove that
-(3.2) $U_{A}(t^{-1})=$ $y^{8}(3^{2}t^{-1}+217\cdot 3+839\cdot 3^{2}t+3140\cdot 3^{2}t^{2}+12070\cdot 3t^{3}-1118\cdot 3^{2}t^{4}$
-$-2126\cdot 3^{3}t^{5}-380\cdot 3^{4}t^{6}+55\cdot 3^{5}t^{7}+19\cdot 3^{6}t^{8}+3^{7}t^{9});$
-(3.3) $U_{A}(1)=$ $y^{8}(t^{-1}+62\cdot 3+100\cdot 3^{3}t+9712t^{2}+3442\cdot 3t^{3}-1588\cdot 3t^{4}-1588\cdot 3^{2}t^{5}$
-$-64\cdot 3^{4}t^{6}+11\cdot 3^{5}t^{7}+2\cdot 3^{6}t^{8});$
-(3.4) $U_{A}(t)=$ $y^{8}(47+328\cdot 3t+1100\cdot 3t^{2}+2792t^{3}-610\cdot 3t^{4}-344\cdot 3^{2}t^{5}-20\cdot 3^{3}t^{6}$
-$+8\cdot 3^{3}t^{7}-3^{4}t^{8});$
-(3.5) $U_{A}(t^{2})=$ $y^{8}(3^{2}+122\cdot 3t+121\cdot 3^{2}t^{2}+29\cdot 3^{3}t^{3}-274\cdot 3t^{4}-40\cdot 3^{2}t^{5}$
-$+22\cdot 3^{3}t^{6}-14\cdot 3^{4}t^{7}-3^{5}t^{8}+2\cdot 3^{6}t^{9}-3^{5}t^{10}-3^{6}t^{11}).$
-Since $L_{1}=U_{A}(L_{0})$ and $L_{0}=t^{-1}+27+3t+9t^{2}$, using (3.2)–(3.5) along with the identities $p_{0}=(1+t)^{4}$, we deduce that
+For example, the expression $L_{0}=t^{-1}+27+3t+9t^{2}$, where $L_{0}$ is defined in Definition 2.3, can be directly verified using this package. The package also yields the identity for $L_{1}$:
 (3.2) $L_{1}=$ $y^{8}p_{0}(4\cdot 3^{2}t^{-1}+71\cdot 3^{4}+2351\cdot 3^{3}t+89\cdot 3^{5}t^{2}-1975\cdot 3^{4}t^{3}+407\cdot 3^{5}t^{4}$
@@
 Definition 3.1.
-With $t=t(\tau)$ as in Definition 2.6 we define:
+Let $t=t(\tau)$ be as in Definition 2.5. We define:
 $a_{0}(t)=-t,~~~a_{1}(t)=3t^{2},~~~a_{2}(t)=-9t^{3}+6t.$
-We define $s:\{0,1,2\}\times\{1,2,3\}\rightarrow\mathbb{Z}$ to be the unique function satisfying
+Define $s:\{0,1,2\}\times\{1,2,3\}\rightarrow\mathbb{Z}$ to be the unique function satisfying
 (3.3) $a_{j}(t)=\sum_{l=1}^{3}s(j,l)3^{\lfloor\frac{2l+j}{3}\rfloor}t^{l}.$
@@
 Proof.
-First we prove
+We first prove that
 $\prod_{\lambda=0}^{2}t_{\lambda}(\tau)=-a_{0}(t)=t.$
-With $\omega:=e^{\frac{2\pi i}{3}}$, we have
+Let $\omega:=e^{\frac{2\pi i}{3}}$. Then
 $\prod_{\lambda=0}^{2}t_{\lambda}(\tau)$ $=\prod_{\lambda=0}^{2}q^{\frac{1}{3}}\omega^{\lambda}\prod_{n=1}^{\infty}\frac{(1-q^{4n})^{4}(1-\omega^{2\lambda n}q^{\frac{2n}{3}})^{2}}{(1-q^{2n})^{2}(1-\omega^{4\lambda n}q^{\frac{4n}{3}})^{4}}$
@@
 $=t,$
-where using the fact that $\prod_{\lambda=0}^{2}(1-\omega^{\lambda n}z)$ equals $(1-z)^{3}$ if $3\mid n$, and $1-z^{3}$ otherwise. For $a_{1}(t)$,
+where we use the fact that $\prod_{\lambda=0}^{2}(1-\omega^{\lambda n}z)$ equals $(1-z)^{3}$ if $3\mid n$, and $1-z^{3}$ otherwise. For $a_{1}(t)$, we have
 $a_{1}(t)=$ $t_{0}(\tau)t_{1}(\tau)+t_{0}(\tau)t_{2}(\tau)+t_{1}(\tau)t_{2}(\tau)=3tU_{3}(t^{-1})=3t^{2},$
-where the last equation was verified by the ETA MAPLE package. For $a_{2}(t)$,
+where the last equation was verified using the ETA MAPLE package. For $a_{2}(t)$, we have
 $a_{2}(t)=$ $-t_{0}(\tau)-t_{1}(\tau)-t_{2}(\tau)=-3U_{3}(t)=-9t^{3}+6t,$
-where the last equation was again verified by the ETA MAPLE package. ∎
-Similarly, we can obtain the following definition and lemma without detailed proofs.
+where the last equation was again verified using the ETA MAPLE package. ∎
+Similarly, we can state the following definition and lemma without detailed proofs.
 Definition 3.3.
@@
 $t_{\lambda}^{3}+a_{2}(t)t_{\lambda}^{2}+a_{1}(t)t_{\lambda}+a_{0}(t)=0.$
-Multiplying both sides with $u_{\lambda}t_{\lambda}^{j}$ where $u_{\lambda}(\tau):=u((\tau+\lambda)/3)$ gives
+Multiplying both sides by $u_{\lambda}t_{\lambda}^{j}$, where $u_{\lambda}(\tau):=u((\tau+\lambda)/3)$, yields
 $u_{\lambda}t_{\lambda}^{j+3}+a_{2}(t)u_{\lambda}t_{\lambda}^{j+2}+a_{1}(t)u_{\lambda}t_{\lambda}^{j+1}+a_{0}(t)u_{\lambda}t_{\lambda}^{j}=0.$
-Summing both sides over all $\lambda$ from $\{0,1,2\}$ completes the proof of (3.8). We can prove (3.9) similarly. ∎
-In the following proofs, we need to use both (3.8) and (3.9), while the coefficients of $y^{m}t^{k}$ are not easily described, where $m\in\mathbb{N}$ and $k\in\mathbb{Z}$.
-According to Group II in the Appendix, it is possible to assume that $U_{A}(p_{1}y^{m}t^{k})$ can be written in the form of $y^{3m+8}p_{0}\sum_{n}a_{n}t^{n}$, $a_{n}\in\mathbb{Z}$ by Lemma 4.3.
-Based on the identity $y=\frac{1}{1-3t}$, we can rewrite $b_{0}(y)$, $b_{1}(y)$, and $b_{2}(y)$ into
+Summing both sides over all $\lambda\in\{0,1,2\}$ completes the proof of (3.4). The identity (3.5) can be proved similarly. ∎
+In the following proofs, we need to use both (3.4) and (3.5).
+However, the coefficients of $y^{m}t^{k}$ are not easily described, where $m\in\mathbb{N}$ and $k\in\mathbb{Z}$.
+According to Group II in the Appendix, by Lemma 4.3, it is possible to assume that $U_{A}(p_{1}y^{m}t^{k})$ can be written in the form $y^{3m+8}p_{0}\sum_{n}a_{n}t^{n}$ with $a_{n}\in\mathbb{Z}$.
+Based on the identity $y=\frac{1}{1-3t}$, we can rewrite $b_{0}(y)$, $b_{1}(y)$, and $b_{2}(y)$ as:
 (3.6) $b_{0}(y)$ $=-y^{9}(1-2\cdot 3^{2}t+5\cdot 3^{3}t^{2}-20\cdot 3^{3}t^{3}+5\cdot 3^{5}t^{4}-2\cdot 3^{6}t^{5}+3^{6}t^{6});$
@@
 (3.8) $b_{2}(y)$ $=-y^{3}(3+4\cdot 3^{2}t+3^{3}t^{2}-2\cdot 3^{3}t^{3}).$
-Similarly, we have $U_{B}(p_{0}y^{m}t^{k})=y^{3m}p_{1}\sum_{n}a_{n}t^{n}$ in Lemma 4.3.
+Similarly, by Lemma 4.3, we have $U_{B}(p_{0}y^{m}t^{k})=y^{3m}p_{1}\sum_{n}a_{n}t^{n}$.
 According to Lemma 3.5, we can calculate $U_{A}(p_{1}y^{m}t^{k})$ and $U_{B}(p_{0}y^{m}t^{k})$ using the identities in Appendix for all $k\geq-1$ and $m\in\mathbb{N}$ recursively.
@@
 Lemma 4.1.
-Let $v,u:\mathbb{H}\rightarrow\mathbb{C}$ and $l\in\mathbb{Z}$. Suppose for $l\leq k\leq l+2$, there exist polynomials $p_{k}(t)\in\mathbb{Z}[t,t^{-1}]$ such that
+Let $v,u:\mathbb{H}\rightarrow\mathbb{C}$ and $\ell\in\mathbb{Z}$. Suppose that for $\ell\leq k\leq\ell+2$, there exist polynomials $p_{k}(t)\in\mathbb{Z}[t,t^{-1}]$ such that
 (4.1) $U_{3}(ut^{k})=vp_{k}(t)$
@@
 (4.2) $ord_{t}(p_{k}(t))\geq\lceil\frac{k+s}{3}\rceil,$
-for a fixed integer $s$. Then there exist families of polynomials $p_{k}(t)\in\mathbb{Z}[t,t^{-1}]$, integer $k\in\mathbb{Z}$, such that (4.1) and (4.2) hold for all $k\in\mathbb{Z}$.
+for some fixed integer $s$.
+Then there exist families of polynomials $p_{k}(t)\in\mathbb{Z}[t,t^{-1}]$, defined for all integers $k\in\mathbb{Z}$, such that both (4.1) and (4.2) hold for all $k\in\mathbb{Z}$.
 Proof.
-Let $N\textgreater l+2$ be an integer and assume by induction that there are families of polynomials $p_{k}(t)$ such that (4.1) and (4.2) hold for $l\leq k\leq N-1$. Suppose
-$p_{k}(t)=\sum_{n\geq\lceil\frac{k+s}{3}\rceil}c(k,n)t^{n},~{}~{}~{}l\leq k\leq N-1,$
-with integers $c(k,n)$. Apply Lemma 3.5 we get:
+Let $N\textgreater\ell+2$ be an integer and assume by induction that there exist families of polynomials $p_{k}(t)$ such that (4.1) and (4.2) hold for all $\ell\leq k\leq N-1$. Suppose
+$p_{k}(t)=\sum_{n\geq\lceil\frac{k+s}{3}\rceil}c(k,n)t^{n},~~~\ell\leq k\leq N-1,$
+with integers $c(k,n)$. Applying Lemma 3.5, we obtain:
 $U_{3}(ut^{N})=$ $-\sum_{j=0}^{2}a_{j}(t)U_{3}(ut^{N+j-3})$
@@
 $=$ $-v\sum_{j=0}^{2}a_{j}(t)t^{-1}\sum_{n\geq\lceil\frac{N+j+s}{3}\rceil}c_{i}(N+j-3,n-1)t^{n}.$
-Recalling the fact that $a_{j}(t)t^{-1}$ for $0\leq j\leq 2$ is a polynomial in $t$, this determines polynomials $p_{N}(t)$ with the desired properties. ∎
+Since $a_{j}(t)t^{-1}$ are polynomials in $t$ for $0\leq j\leq 2$, this defines a polynomial $p_{N}(t)$ with the desired properties. The induction argument for $N\textless\ell$ is analogous. ∎
 Lemma 4.2.
-Let $v,u:\mathbb{H}\rightarrow\mathbb{C}$ and $l\in\mathbb{Z}$. Suppose for $l\leq k\leq l+2$ there exist polynomials $p_{k}(t)\in\mathbb{Z}[t,t^{-1}]$, such that
+Let $v,u:\mathbb{H}\rightarrow\mathbb{C}$ and $\ell\in\mathbb{Z}$. Suppose that for $\ell\leq k\leq\ell+2$ there exist polynomials $p_{k}(t)\in\mathbb{Z}[t,t^{-1}]$, such that
 (4.3) $U_{3}(ut^{k})=vp_{k}(t)$
-where
+with
 (4.4) $p_{k}(t)=\sum_{n}c(k,n)3^{\lfloor\frac{2n-k+\gamma}{3}\rfloor}t^{n},$
-with integers $\gamma$ and $c(k,n)$.
-Then there exist families of polynomials $p_{k}(t)\in\mathbb{Z}[t,t^{-1}]$, $k\in\mathbb{Z}$, of the form (4.4) for which property (4.3) holds for all $k\in\mathbb{Z}$.
+for some integers $\gamma$ and $c(k,n)$.
+Then there exist families of polynomials $p_{k}(t)\in\mathbb{Z}[t,t^{-1}]$, defined for all $k\in\mathbb{Z}$, of the form (4.4) such that property (4.3) holds for all $k\in\mathbb{Z}$.
 Proof.
-Suppose for an integer $N\textgreater l+2$ there are families of polynomials $p_{k}(t)$, of the form (4.4) satisfying property (4.3) for $l\leq k\leq N-1$.
+Suppose for an integer $N\textgreater\ell+2$ there exist families of polynomials $p_{k}(t)$ of the form (4.4) satisfying property (4.3) for all $\ell\leq k\leq N-1$.
 We proceed by mathematical induction on $N$.
-Applying Lemma 3.5 and using the induction base (4.3) and (4.4) we attain:
+Applying Lemma 3.5 and using the induction hypothesis (4.3) and (4.4), we obtain:
 $U_{3}(ut^{N})=-\sum_{j=0}^{2}a_{j}(t)v\sum_{n}c(N+j-3,n)3^{\lfloor{\frac{2n-(N+j-3)+\gamma}{3}}\rfloor}t^{n}.$
-Utilizing (3.7) this rewrites into:
-$U_{3}(ut^{N})$ $=-\sum_{j=0}^{2}\sum_{l=1}^{3}s(j,l)3^{\lfloor{\frac{2l+j}{3}}\rfloor}t^{l}v\sum_{n}c(N+j-3,n)3^{\lfloor{\frac{2n-(N+j-3)+\gamma}{3}}\rfloor}t^{n}$
-$=-v\sum_{j=0}^{2}\sum_{l=1}^{3}\sum_{n}s(j,l)c(N+j-3,n-l)3^{\lfloor{\frac{2(n-l)-(N+j-3)+\gamma}{3}}\rfloor+\lfloor{\frac{2l+j}{3}}\rfloor}t^{n}.$
+Utilizing (3.3), this becomes:
+$U_{3}(ut^{N})$ $=-\sum_{j=0}^{2}\sum_{\ell=1}^{3}s(j,l)3^{\lfloor{\frac{2\ell+j}{3}}\rfloor}t^{\ell}v\sum_{n}c(N+j-3,n)3^{\lfloor{\frac{2n-(N+j-3)+\gamma}{3}}\rfloor}t^{n}$
+$=-v\sum_{j=0}^{2}\sum_{l=1}^{3}\sum_{n}s(j,\ell)c(N+j-3,n-\ell)3^{\lfloor{\frac{2(n-\ell)-(N+j-3)+\gamma}{3}}\rfloor+\lfloor{\frac{2\ell+j}{3}}\rfloor}t^{n}.$
 The induction step is completed by simplifying the exponent of $3$ as follows:
-$\lfloor{\frac{2(n-l)-(N+j-3)+\gamma}{3}}\rfloor+\lfloor{\frac{2l+j}{3}}\rfloor$
-$\geq\lfloor{\frac{2(n-l)-(N+j-3)+\gamma+2l+j-2}{3}}\rfloor$
+$\lfloor{\frac{2(n-\ell)-(N+j-3)+\gamma}{3}}\rfloor+\lfloor{\frac{2\ell+j}{3}}\rfloor$
+$\geq\lfloor{\frac{2(n-\ell)-(N+j-3)+\gamma+2\ell+j-2}{3}}\rfloor$
 $\geq\lfloor{\frac{2n-N+\gamma}{3}}\rfloor.$
-The induction proof for $N\textless l$ works analogously. ∎
+The induction argument for $N\textless\ell$ proceeds analogously. ∎
 Lemma 4.3.
-Let $v,u:\mathbb{H}\rightarrow\mathbb{C}$, $l\in\mathbb{Z}$, and $r\in\mathbb{N}$.
-Suppose for $l\leq k\leq l+2$ and $r\leq m\leq r+2$ there exist polynomials $p_{m,k}(t)\in\mathbb{Z}[t,t^{-1}]$ such that
+Let $v,u:\mathbb{H}\rightarrow\mathbb{C}$, $\ell\in\mathbb{Z}$, and $r\in\mathbb{N}$.
+Suppose that for $\ell\leq k\leq\ell+2$ and $r\leq m\leq r+2$, there exist polynomials $p_{m,k}(t)\in\mathbb{Z}[t,t^{-1}]$ such that
 (4.5) $U_{3}(uy^{m}t^{k})=vy^{3m+a}p_{m,k}(t)$
@@
 for some fixed integers $a$ and $s$.
-Then there exist families of polynomials $p_{m,k}(t)\in\mathbb{Z}[t,t^{-1}]$, integer $k\in\mathbb{Z}$, such that (4.5) and (4.6) hold for all $m\in\mathbb{Z^{\ast}}$ and $k\in\mathbb{Z}$.
+Then there exist families of polynomials $p_{m,k}(t)\in\mathbb{Z}[t,t^{-1}]$, defined for all integers $k\in\mathbb{Z}$, such that both (4.5) and (4.6) hold for all $m\geq r$ and $k\in\mathbb{Z}$.
 Proof.
-By Lemma 4.1, we can directly get that for $r\leq m\leq r+2$, (4.5) and (4.6) hold for all $k\in\mathbb{Z}$.
-For a fixed integer $k$, let $N\textgreater r+2$ be an integer and assume by induction that there are families of polynomials $p_{m,k}(t)$ such that (4.5) and (4.6) hold for $r\leq m\leq N-1$.
+By Lemma 4.1, we conclude that for $r\leq m\leq r+2$, both (4.5) and (4.6) hold for all $k\in\mathbb{Z}$.
+Now fix an integer $k$, and let $N\textgreater r+2$ be an integer.
+Assume by induction that there exist families of polynomials $p_{m,k}(t)$ such that (4.5) and (4.6) hold for all $r\leq m\leq N-1$.
 Suppose
 $p_{m,k}(t)=\sum_{n\geq\lceil\frac{k+s}{3}\rceil}c(k,m,n)t^{n},~~~r\leq m\leq N-1,$
-with integers $c(k,m,n)$. Apply Lemma 3.5 we get:
+with integers $c(k,m,n)$. Applying Lemma 3.5, we obtain:
 $U_{3}(uy^{N}t^{k})=$ $-\sum_{j=0}^{2}b_{j}(y)U_{3}(uy^{N+j-3}t^{k})$
 $=$ $-\sum_{j=0}^{2}b_{j}(y)vy^{3(N+j-3)+a}\sum_{n\geq\lceil\frac{k+s}{3}\rceil}c(k,m,n)t^{n}.$
-Recalling the equations (3.10)–(3.12) this determines (4.5) and (4.6) hold for all $m\in\mathbb{N}$. Thus we complete the proof of this lemma. ∎
+Using equations (3.6)–(3.8), we verify that both (4.5) and (4.6) hold for $m=N$. By induction, it follows that these properties hold for all $m\geq r$, which completes the proof. ∎
 Lemma 4.4.
-Let $v,u:\mathbb{H}\rightarrow\mathbb{C}$, $l\in\mathbb{Z}$, and $r\in\mathbb{N}$.
-Suppose for $l\leq k\leq l+2$ and $r\leq m\leq r+2$ there exist polynomials $p_{m,k}(t)\in\mathbb{Z}[t,t^{-1}]$ such that
+Let $v,u:\mathbb{H}\rightarrow\mathbb{C}$, $\ell\in\mathbb{Z}$, and $r\in\mathbb{N}$.
+Suppose that for $\ell\leq k\leq\ell+2$ and $r\leq m\leq r+2$, there exist polynomials $p_{m,k}(t)\in\mathbb{Z}[t,t^{-1}]$ such that
 (4.7) $U_{3}(uy^{m}t^{k})=vy^{3m+a}p_{m,k}(t)$
@@
 (4.8) $p_{m,k}(t)=\sum_{n}c(k,m,n)3^{\lfloor\frac{2n-k+\gamma}{3}\rfloor}t^{n},$
-with integers $\gamma$ and $c(k,m,n)$.
-Then there exist families of polynomials $p_{m,k}(t)\in\mathbb{Z}[t,t^{-1}]$, $k\in\mathbb{Z}$, of the form (4.8) for which property (4.7) holds for all $m\in\mathbb{N}$ and $k\in\mathbb{Z}$.
+for some integers $\gamma$ and $c(k,m,n)$.
+Then there exist families of polynomials $p_{m,k}(t)\in\mathbb{Z}[t,t^{-1}]$, defined for all $k\in\mathbb{Z}$ and of the form (4.8), such that property (4.7) holds for all $m\geq r$ and $k\in\mathbb{Z}$.
 Proof.
-By Lemma 4.2, we can directly get that for $r\leq m\leq r+2$, (4.7) and (4.8) hold for all $k\in\mathbb{Z}$.
-For a fixed integer $k$, let $N\textgreater r+2$ be an integer and assume by induction that there are families of polynomials $p_{m,k}(t)$ such that (4.7) and (4.8) hold for $r\leq m\leq N-1$.
+By Lemma 4.2, we conclude that for $r\leq m\leq r+2$, both (4.7) and (4.8) hold for all $k\in\mathbb{Z}$.
+Now fix an integer $k$, and let $N\textgreater r+2$ be an integer.
+Assume by induction that there exist families of polynomials $p_{m,k}(t)$ such that (4.7) and (4.8) hold for all $r\leq m\leq N-1$.
 We proceed by mathematical induction on N.
-Applying Lemma 3.5 and using the induction base (4.7) and (4.8) we attain:
+Applying Lemma 3.5 and using the induction hypothesis (4.7) and (4.8), we attain:
 $U_{3}(uy^{N}t^{k})=-\sum_{j=0}^{2}b_{j}(y)U_{3}(uy^{N+j-3}t^{k})=-v\sum_{j=0}^{2}b_{j}(y)y^{3(N+j-3)+a}\sum_{n}c(k,m,n)3^{\lfloor\frac{2n-k+\gamma}{3}\rfloor}t^{n}.$
-Since $b_{j}(y)$ can be written into $y^{9-3j}f(t)$, where $j\in\{0,1,2\}$ and $ord_{t}(f(t))\geq 0$, then (4.7) and (4.8) hold for $m=N$. Thus we complete the proof of this lemma. ∎
+Since each $b_{j}(y)$ can be expressed as $y^{9-3j}f(t)$ for $j\in\{0,1,2\}$, where $ord_{t}(f(t))\geq 0$, it follows that both (4.7) and (4.8) hold for $m=N$.
+By in duction, this completes the proof of this lemma.
+∎
 Definition 4.5.
 A map $a:\mathbb{Z}\rightarrow\mathbb{Z}$ is called a discrete function if it has finite support.
-A map $a:\mathbb{Z}\times\mathbb{N}\times\mathbb{Z}\rightarrow\mathbb{Z}$ is called discrete array if for each $i\in\mathbb{Z}$, $j\in\mathbb{N}$ the map $a(i,j,-):\mathbb{Z}\rightarrow\mathbb{Z},k\mapsto a(i,j,k)$, has finite support.
+A map $a:\mathbb{Z}\times\mathbb{N}\times\mathbb{Z}\rightarrow\mathbb{Z}$ is called a discrete array if for each $i\in\mathbb{Z}$ and $j\in\mathbb{N}$, the map $a(i,j,-):\mathbb{Z}\rightarrow\mathbb{Z},k\mapsto a(i,j,k)$ has finite support.
 Lemma 4.6.
-Given $U_{A}(f)$, $U_{B}(f)$ as in Definition 2.2, $p_{0},p_{1}$, and $y$ as in Definition 2.6.
-Then there exist discrete arrays $a$ and $b$, such that the following relations hold for all $m\in\mathbb{N}$, $k\in\mathbb{Z}$:
+Let $U_{A}(f)$, $U_{B}(f)$ be as in Definition 2.2, and let $p_{0},p_{1}$, and $y$ be as in Definition 2.5.
+Then there exist discrete arrays $a$ and $b$ such that the following relations hold for all $m\in\mathbb{N}$ and $k\in\mathbb{Z}$:
 (4.9) $U_{A}(p_{1}y^{m}t^{k})=y^{3m+8}p_{0}\sum_{n\geq\lceil\frac{k-3}{3}\rceil}a(k,m,n)t^{n},\text{where}~\pi(a(k,m,n))\geq\small{\lfloor\frac{2n-k+3}{3}\rfloor};$
@@
 The Appendix lists eighteen fundamental relations.
-The nine fundamental relations of Group II fit the pattern of the relation (4.9) for three values of $m$ and three values of $k$.
-The same observations apply to the Groups III with regard to the relations (4.10).
-Applying Lemmas 4.14.4, we immediately prove the statement for all $m\in\mathbb{N}$ and $k\in\mathbb{Z}$.
+The nine fundamental relations in Group II match the pattern of the relation (4.9) for three values of $m$ and three values of $k$.
+Similarly, the relations in Groups III correspond to the pattern of (4.10).
+Applying Lemmas 4.14.4, we immediately obtin the statement for all $m\in\mathbb{N}$ and $k\in\mathbb{Z}$.
 ∎
-In the Appendix, we mention the identity that $y=\frac{1}{1-3t}$.
-In the following proof, we will express each $L_{\alpha}$ as a rational polynomial in which the denominator consists of powers of $1-3t$.
-With in this mind, we define the multiplicative closed set
+In the following proof, we use the next lemma to describe the powers of $y$ in $L_{\alpha}$.
+Lemma 4.7.
+Define $f(1)=8$, and for $s\geq 2$, dedine $f(s)$ recursively by:
+$f(s)=\left\{\begin{array}[]{ll}3f(s-1),&\text{if}~s~\text{is even,}\\ 3f(s-1)+8,&\text{if}~s~\text{is odd.}\end{array}\right$.
+Then for all $s\geq 1$,
+$f(s)=\left\{\begin{array}[]{ll}3^{s+1}-3,&\text{if}~s~\text{is even,}\\ 3^{s+1}-1,&\text{if}~s~\text{is odd.}\end{array}\right$.
+In the Appendix, we mention the identity $y=\frac{1}{1-3t}$.
+In the following proof, we will express each $L_{\alpha}$ as a rational polynomial in which the denominator is a power of 13t1-3t.
+With this in mind, we define the multiplicative closed set
 $\mathcal{S}:=\{(1-3t)^{n}:n\in\mathbb{Z}_{n\geq 0}\}.$
-We will be working with subspaces of the localized ring
+We will work with subspaces of the localized ring
 $\mathbb{Z}[t,t^{-1}]_{\mathcal{S}}:=\mathcal{S}^{-1}\mathbb{Z}[t,t^{-1}].$
-Given the form of $L_{1}$ in (3.6), we expect to represent each $L_{\alpha}$ in terms of $\frac{t^{n}}{(1-3t)^{m}}$ for $n\geq-1$ and $m\geq 0$. If we set
-$L_{2\alpha-1}=p_{0}y^{3^{2\alpha}-1}\sum_{n\geq-1}d_{n}^{(2\alpha-1)}t^{n};$
+Given the form of $L_{1}$ in (3.2), we expect to represent each $L_{\alpha}$ in the form $\frac{t^{n}}{(1-3t)^{m}}$ for $n\geq-1$ and $m\geq 0$. If we set
+$L_{2\alpha-1}=p_{0}y^{3^{2\alpha}-1}\sum_{n\geq-1}d_{n}^{(2\alpha-1)}t^{n},$
 $L_{2\alpha}=p_{1}y^{3^{2\alpha+1}-3}\sum_{n\geq 0}d_{n}^{(2\alpha)}t^{n},$
-for $\alpha\geq 1$. Recalling that $\pi(n)$ denotes the 3-adic order of n (i.e. the highest power of 3 that divides $n$), then Theorem 1.1 follows from the next lemma.
+for $\alpha\geq 1$, and recall that $\pi(n)$ denotes the 3-adic order of n (i.e. the highest power of 3 that divides $n$), then Theorem 1.1 follows from the next lemma.
 Lemma 4.8.
-For integer $\alpha\geq 1$, we have
+For every integer $\alpha\geq 1$, we have
 (4.11) $\pi(d_{n}^{(2\alpha-1)})\geq\lfloor\frac{2n+5}{3}\rfloor+\alpha,~~\text{for $n\geq-1$},$
@@
 $-19\cdot 3^{7}t^{5}+11\cdot 3^{7}t^{6}-3^{8}t^{7}),$
-which demonstrates that (4.11) and (4.13) hold for $\alpha=1$. Next, we suppose (4.11) and (4.13) hold for $\alpha\leq s$. Then by Definition 2.4 and Lemma 4.6, we have
+which shows that $\pi(d_{n}^{(1)})\geq\lfloor\frac{2n+5}{3}\rfloor+1$ and $\pi(d_{0}^{(1)})=4$. Thus, (4.11) and (4.13) hold for $\alpha=1$.
+Next, suppose (4.11) and (4.13) hold for $\alpha\leq s$.
+We want to use $U_{B}$ operator on $L_{2s-1}$ to get $L_{2s}$, especially the coefficients $d_{n}^{(2s)}$ for $n\geq 0$.
+Then by Definition 2.3 and Lemma 4.6, we have
 $L_{2s}=$ $U_{B}(L_{2s-1})$
@@
 $=$ $p_{1}y^{3^{2s+1}-3}\sum_{n\geq 0}\sum_{-1\leq k\leq 3n+1}d_{k}^{(2s-1)}b(k,3^{2s}-1,n)t^{n}.$
-By comparing the coefficients of $p_{1}y^{3^{2s+1}-3}t^{n}$ in the equation above, we have
+Comparing the coefficients of $p_{1}y^{3^{2s+1}-3}t^{n}$ on both sides, we have
 (4.15) $d_{n}^{(2s)}=\sum_{-1\leq k\leq 3n+1}d_{k}^{(2s-1)}b(k,3^{2s}-1,n).$
-Hence, for $n\geq 1$, we have
-$\pi(d_{n}^{(2s)})$ $\geq\min_{k\geq-1}\{\lfloor\frac{2k+5}{3}\rfloor+s+\lfloor\frac{2n-k+2}{3}\rfloor\}$
+We notice that $\pi(n)$ has the properties that $\pi(x+y)\geq\min\{\pi(x),\pi(y)\}$ and $\pi(xy)=\pi(x)+\pi(y)$ for $x,y\in\mathbb{Z}$.
+According to (4.10) and (4.11), we obtain that $\pi(b(k,3^{2s}-1,n))\geq\lfloor\frac{2n-k+2}{3}\rfloor$ and $\pi(d_{k}^{(2s-1)})\geq\lfloor\frac{2k+5}{3}\rfloor+s$.
+Hence, for $n\geq 1$,
+$\pi(d_{n}^{(2s)})$ $\geq\min_{k\geq-1}\{\pi(d_{k}^{(2s-1)})+\pi(b(k,3^{2s}-1,n))\}$
+$\geq\min_{k\geq-1}\{\lfloor\frac{2k+5}{3}\rfloor+s+\lfloor\frac{2n-k+2}{3}\rfloor\}$
 $\geq\lfloor\frac{2n+5}{3}\rfloor+s,$
-which indicates (4.12) holds for $\alpha=s$. For $d_{0}^{(2s)}$, from (4.15) we obtain that
+which indicates that (4.12) holds for $\alpha=s$.
+For $d_{0}^{(2s)}$, from (4.15), we obtain that
 $d_{0}^{(2s)}=$ $d_{-1}^{(2s-1)}b(-1,3^{2s}-1,0)+d_{0}^{(2s-1)}b(0,3^{2s}-1,0)+d_{1}^{(2s-1)}b(1,3^{2s}-1,0).$
-By the assumption, we have $\pi(d_{-1}^{(2s-1)})\geq s+1$, $\pi(d_{0}^{(2s-1)})\geq s+2$, and $\pi(d_{1}^{(2s-1)})\geq s+2$.
-We denote $[t^{k}]f$ to the coefficients of $t^{k}$ in $f$.
-With (3.9), we have
+By the induction hypothesis, $\pi(d_{-1}^{(2s-1)})\geq s+1$, $\pi(d_{0}^{(2s-1)})\geq s+2$, and $\pi(d_{1}^{(2s-1)})\geq s+2$.
+Let $t^{k}]f$ denote the coefficients of $t^{k}$ in $f$. Using(3.5), we find
 (4.16) $b(-1,m,0)=$ $[p_{1}y^{3m}]U_{B}(p_{0}y^{m}t^{-1})$
@@
 $=$ $b(-1,m-3,0)-3b(-1,m-2,0)+3b(-1,m-1,0).$
-In the Appendix, we get that $b(-1,0,0)=12$, $b(-1,1,0)=36$, and $b(-1,2,0)=69$ which all can be divided by $3$.
-By the recurrence relation above, we attain that $3\mid b(-1,m,0)$ for all $m\in\mathbb{N}$.
-Thus, we get that (4.14) holds for $\alpha=s$.
-Finally, we suppose (4.12) and (4.14) hold for $\alpha\leq s$. Then by Definition 2.4 and Lemma 4.6, we have
+From the Appendix, b(1,0,0)=12b(-1,0,0)=12, b(1,1,0)=36b(-1,1,0)=36, and b(1,2,0)=69b(-1,2,0)=69, all divisible by $3$.
+By the recurrence above, $3\mid b(-1,m,0)$ for all $m\in\mathbb{N}$.
+Hence, (4.14) holds for $\alpha=s$.
+Now suppose (4.12) and (4.14) hold for $\alpha\leq s$. Then by Definition 2.3 and Lemma 4.6,
 $L_{2s+1}=$ $U_{A}(L_{2s})$
@@
 $=$ $p_{0}y^{3^{2s+2}-1}\sum_{n\geq-1}\sum_{0\leq k\leq 3n+3}d_{k}^{(2s)}a(k,3^{2s+1}-3,n)t^{n}.$
-Equating the coefficients of $p_{0}y^{3^{2s+2}-1}t^{n}$ in the equation above, we have
+Equating coefficients of $p_{0}y^{3^{2s+2}-1}t^{n}$ on both sides, we get
 (4.17) $d_{n}^{(2s+1)}=\sum_{0\leq k\leq 3n+3}d_{k}^{(2s)}a(k,3^{2s+1}-3,n).$
-Hence, for $n\geq-1$, we have
-$\pi(d_{n}^{(2s+1)})$ $\geq\min\{\min_{k\geq 1}\{\lfloor\frac{2k+5}{3}\rfloor+s+\lfloor\frac{2n-k+3}{3}\rfloor\},s+2+\lfloor\frac{2n+3}{3}\rfloor\}$
+According to (4.9), (4.12), and (4.14), we obtain that $\pi(a(k,3^{2s+1}-3,n))\geq\lfloor\frac{2n-k+3}{2}\rfloor$, $\pi(d_{k}^{(2s)})\geq\lfloor\frac{2n+5}{3}\rfloor+s$ for $k\geq 1$, and $\pi(d_{0}^{(2s)})\geq s+2$.
+Thus, for $n\geq-1$,
+$\pi(d_{n}^{(2s+1)})$ $\geq\min_{k\geq 0}\{\pi(d_{k}^{(2s)})+\pi(a(k,3^{2s+1}-3,n))\}$
+$\geq\min\{\min_{k\geq 1}\{\pi(d_{k}^{(2s)})+\pi(a(k,3^{2s+1}-3,n))\},\pi(d_{0}^{(2s)})+\pi(a(0,3^{2s+1}-3,n))\}$
+$\geq\min\{\min_{k\geq 1}\{\lfloor\frac{2k+5}{3}\rfloor+s+\lfloor\frac{2n-k+3}{3}\rfloor\},s+2+\lfloor\frac{2n+3}{3}\rfloor\}$
+$\geq\min\{\min\{\min_{k\geq 2}\{\lfloor\frac{2n+k+3}{3}\rfloor+s+1\},\lfloor\frac{2n+5}{3}\rfloor+s+1\},\lfloor\frac{2n+6}{3}\rfloor+s+1\}$
 $\geq\lfloor\frac{2n+5}{3}\rfloor+s+1,$
-which indicates (4.11) holds for $\alpha=s+1$. From (4.17), we have
+where the forth inequality is obtained by splitting the part where $k\textgreater 1$ into two cases: $k=1$ and $k\geq 2$. Thus, (4.11) holds for $\alpha=s+1$.
+From (4.17), we have
 $d_{0}^{(2s+1)}=$ $d_{0}^{(2s)}a(0,3^{2s+1}-3,0)+d_{1}^{(2s)}a(1,3^{2s+1}-3,0)+d_{2}^{(2s)}a(2,3^{2s+1}-3,0)$
 $+d_{3}^{(2s)}a(3,3^{2s+1}-3,0).$
-By the assumption, we have $\pi(d_{0}^{(2s)})\geq s+2$, $\pi(d_{1}^{(2s)})\geq s+2$, $\pi(d_{2}^{(2s)})\geq s+3$, and $\pi(d_{3}^{(2s)})\geq s+3$. Similar to (4.16), we obtain that
+By the induction hypothesis, $\pi(d_{0}^{(2s)})\geq s+2$, $\pi(d_{1}^{(2s)})\geq s+2$, $\pi(d_{2}^{(2s)})\geq s+3$, and $\pi(d_{3}^{(2s)})\geq s+3$.
+Similar to (4.16), we derive
 $a(0,m,0)=$ $a(0,m-3,0)-18a(0,m-3,-1)-3a(0,m-2,0)+9a(0,m-2,-1)$
@@
 $a(1,m,0)=$ $a(1,m-3,0)-3a(1,m-2,0)+3a(1,m-1,0).$
-From the Appendix, we get that
+From the Appendix,
 $a(0,0,0)=285,\hskip 10.0pta(0,1,0)=600,\hskip 10.0pta(0,2,0)=1068;$
 $a(1,0,0)=66,\hskip 16.0pta(1,1,0)=108,\hskip 10.0pta(1,2,0)=159.$
-Hence, we attain that $3\mid a(0,m,0)$ and $3\mid a(1,m,0)$ for all $m\in\mathbb{N}$ so that (4.13) holds for $\alpha=s+1$, which completes the proof of Lemma 4.7. ∎
+Hence, $3\mid a(0,m,0)$ and $3\mid a(1,m,0)$ for all $m\in\mathbb{N}$, so (4.13) holds for $\alpha=s+1$. This completes the proof of Lemma 4.8. ∎
 Appendix. The fundamental relations
@@
 Acknowledgements
-The first author was supported by the National Key R&D Program of China (Grant No.
-2024YFA1014500).
-The first author was also supported in part by the National Natural Science Foundation of China (Grant No.
-12201387).
+The first author was supported by the National Key R&D Program of China (Grant No. 2024YFA1014500) and the National Natural Science Foundation of China (Grant No. 12201387).